### v2

# Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology

Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the Reverse Cuthill McKee Algorithm (RCM) algorithm is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide orthogonal blocks into one category, which greatly reduces computational complexity. In addition, the convergence of POBK algorithm is theoretically proven, and a corresponding analysis of its faster convergence is also provided. Finally, numerical simulations indicate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k) and GREBK(k) in both iteration steps (IT) and CPU time aspects.
keywords:
Kaczmarz algorithm, Preprocessing techniques, Orthogonal partitioning, Robustness, k-means
journal:
1 Introduction
This paper mainly considers the solution of a large uniformly sparse linear system in the form of
$Ax=f$ (1.1)
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. One of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method.
The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i}=\{x\mid\left\langle A_{i},x\right\rangle=f_{i}\}$ defined by the system rows 1 ; 2 ; 19 . Specifically, the initial value is set to $x_{0}$, the iteration format of the classic Kaczmarz method is
$x_{j+1}=x_{j}+\frac{f_{i}-A_{i}^{T}x}{\left\|A_{i}\right\|_{2}^{2}}\left(A_{i}^{T}\right).$ (1.2)
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin 3 ; 4 proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method 5 . Nutini et al. proposed the greedy distance Kaczmarz (GDK) method 6 . Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version, which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem 9 . Du Yishu et al. introduced an effective distance based row selection criterion and proposed the distance based greedy random Kaczmarz (GDRK) algorithm 10 , which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method 11 ; 12 is a natural extension of the classical Kaczmarz method and another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row metrics in each iteration. Given an initial solution $x_{0}$, in the k-th iteration, first select a subset of travel indicators $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
$x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right),$ (1.3)
where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$. Needell et al. proposed the Good Saving and Block Random Kaczmarz (RBK) algorithms 13 ; 14 ; 15 . Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) 18 . Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual 16 . Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK (k)) algorithm based on the k-means algorithm 17 and the global random block Kaczmarz method 18 , further improving the convergence rate of the algorithm. On this basis, Zheng Wenhao et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK (k)) algorithm 20 , which has a much faster computational speed than the RBK (k) method.
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. When faced with sparsely distributed matrices, the iteration time for the block Kaczmarz algorithm to obtain accurate solutions is not ideal. To improve the convergence speed of block Kaczmarz and the convergence of iterative sequences, this paper constructs an orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques. The algorithm mainly uses preprocessing techniques to rearrange the original matrix, making the new matrix an approximate block diagonal matrix. Then, the matrix is divided into blocks, with the aim of grouping orthogonal blocks into one category. In this way, each iteration in each category undergoes two orthogonal projections on two mutually orthogonal hyperplanes, so the iterative solution can approach the exact solution more quickly. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), and GREBK(k) algorithms, the POBK algorithm is more efficient and ideal in terms of iterative sequence convergence and algorithm running time.
The structure of this article is as follows. In Section 2, we briefly introduced the GRBK algorithm, RBK (k) algorithm, and GREBK (k) algorithm. In Section 3, we introduced the preprocessing techniques and partitioning ideas used, analyzed the reasons for the selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected appropriate matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm 17
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Randomly block $A$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ for row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK (k)) method 17 . Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block sub-matrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
Algorithm 2 RBK(k) Algorithm 17
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $A$ and $f$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2023, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals 20 , that is, the standardized residuals are the 2-norm of the row vector of the ordinary residuals divided by the corresponding sparse matrix $A$
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{i}^{T}\right\|_{2}}=\frac{f_{i}-A_{i}^{T}x_{0}}{\left\|A_{i}^{T}\right\|_{2}}.$ (2.1)
When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
$d_{i}=\lvert{\hat{r}_{i}}\rvert$ (2.2)
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the blocks of $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to block the standardized residuals clustering, as shown in Algorithm 3.
Algorithm 3 GREBK(k) Algorithm 20
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
3 Orthogonal Blocking Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Reverse Cuthill McKee Algorithm
As is well known, the Kaczmarz algorithm has good convergence for linear systems, especially for sparse linear systems. However, Kaczmarz algorithm does not exhibit good convergence speed for all sparse matrices, especially for matrices with large dimensions or scattered structural distributions. Even using block clustering and greedy algorithms cannot effectively change this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the structure of matrices in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed and the angle between the hyperplane where the iteration point is located and the next projection hyperplane are closely related. Obviously, the smaller the angle between the two hyperplanes, the more iterations, and the slower the convergence. The closer the angle is to the perpendicular, the fewer iterations, and the faster the convergence. Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$, It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see 21 .
Figure 1:  Comparison of Kaczmarz Iterations on Small and Vertical Angles on a 2D System.
The Reverse Cuthill McKee (RCM) algorithm was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., $\widetilde{A}$ has diagonal elements closer to the diagonal.
Fortunately, the RCM algorithm also performs equally well on asymmetric matrices. So, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm, transforming the original coefficient starting matrix into an approximate (block) diagonal matrix, see Figure 2. Obviously, there are many rows or row blocks that are mutually orthogonal in the approximate (block) diagonal matrix, see Figure 3. And this improves the drawback of the original matrix structure being arranged randomly. According to our above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). Taking matrices Poli3, Blackhole, Dixmaanl, and Jagmesh4 as examples, we visualized the sparse structure of these matrices and the matrices that were reordered by the RCM algorithm. We compared the structural changes before and after processing with the RCM algorithm, as shown in Figure 2. The blue part represents the distribution of non-zero elements in the matrix. From Figure 2, it is evident that after being rearranged by the RCM algorithm, the non-zero elements of the matrix are closer to the diagonal. Therefore, choosing the RCM algorithm for preprocessing is reasonable and effective.
Figure 2:   The top four images show the sparse structure of matrices Poli3, Blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
3.2 Proposed methods
As mentioned above, both RBK(k) and GREK(k) algorithms use the k-means algorithm to group matrix rows. Undoubtedly, k-means clustering is one of the most effective methods for dividing data into different groups. However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high. Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$, and the preprocessed matrix is
$PAP^{T}$ (3.1)
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements in the original coefficient matrix $A$ closer to the main diagonal elements. Even if the matrix structure is closer to block diagonal form and $P$ is a permutation matrix (i.e., with $P^{T}P=E$), the linear system (1.1) becomes
$PAP^{T}Px=Pf.$ (3.2)
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate block diagonal matrix, as long as $\widetilde{A}$ is appropriately partitioned, orthogonality between blocks can be achieved, see Figure 3.
Figure 3:   The orthogonality representation of the matrix Blackhole after being rearranged using the RCM algorithm.
Consider dividing $\widetilde{A}$ into $k$ blocks, taking $\left[m/k\right]$ elements from each block, and storing the remaining elements in the $k+1$-th block. Perform corresponding partitioning on $\widetilde{f}$, that is, partitioning:
$\widetilde{A}=\begin{bmatrix}A_{\tau_{1}}\\ A_{\tau_{2}}\\ ...\\ A_{\tau_{k}}\\ A_{\tau_{k+1}}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{\tau_{1}}\\ f_{\tau_{2}}\\ ...\\ f_{\tau_{k}}\\ f_{\tau_{k+1}}\\ \end{bmatrix}.$
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. So obviously, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non orthogonal hyperplanes. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{\tau_{i}}$ as $\overline{A}_{\tau_{i}}$, $i=1,2,...,k+1$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates to form an inner product table $G$, $G\left(i,j\right)$ which is used to measure the orthogonality between blocks $\overline{A}_{\tau_{i}}$ and $\overline{A}_{\tau_{j}}$. Moreover, as $\widetilde{A}$ is an approximate block diagonal matrix, it is unreasonable and convenient to use the inner product of centroid coordinates to reflect the orthogonality between blocks. And we divide the matrix $\widetilde{A}$ into $k+1$ blocks in advance, in fact, we only need to do inner product on the centroid coordinates of these $k+1$ blocks. Moreover, since $G$ is a symmetric matrix, we only need to calculate $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times, because $k\ll m$, this is a relatively small number.
Next, based on our above analysis, our aim is to group two row blocks that are mutually orthogonal, i.e., $G\left(i,j\right)=0$ into one class to accelerate the Kaczmarz method. However, it is evident that there may be more than one block orthogonal to any block, which can easily result in some row blocks being in multiple categories simultaneously. In order to avoid this situation, we stipulate that for each block $A_{\tau_{i}}$, $i=1,2,..k$, we select the first block in $\left\{A_{\tau_{i}},A_{\tau_{i+1}},...,A_{\tau_{k+1}}\right\}$ that has an inner product of 00 with the centroid coordinates of the current block $A_{\tau_{i}}$ and place it in one category, marking the current class as $class\left\{{i}\right\}$, and ensuring that the blocks between each class are not duplicated. For blocks that have not entered any class, in order to avoid non-convergence caused by information loss, we classify all these blocks into a new class, denoted as $class\left\{l\right\}$. This regulation is reasonable because $\widetilde{A}$ approximates the properties of block diagonal matrices. Obtain solutions through block Kaczmarz iteration method on two categories
$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
performing two iterations on pairwise orthogonal blocks greatly accelerates the convergence speed of the Kaczmarz algorithm, and then performing Kaczmarz iterations on non-orthogonal blocks ensures that matrix information is not omitted.
Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$.
Algorithm 4 POBK Algorithm
1:sparse matrix $A$, right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, the maximum number of iterations $l$.
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
4:Apply the k-means method to $\widetilde{A}$ to obtain a partition $\{\tau_{1},...,\tau_{k+1}\}$ of the row $\{1,...m\}$, and obtain $k+1$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k+1$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
5:Calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates of each block to form the inner product table $G$.
6:According to the above regulations, place mutually orthogonal blocks in $class\left\{i\right\}$, and ensure that the blocks between each class are not duplicated. Divide all remaining blocks into a new $class\left\{l\right\}$.
7:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
8:     Select   $\tau_{j1},\tau_{j2}\in class\left\{{i}\right\}$
9:     Compute
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right)$
$\widetilde{y}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{j2}}^{\dagger}\left(f_{\tau_{j2}}-A_{\tau_{j2}}\widetilde{x}_{{j+1}/2}\right)$
10:     Select   $\tau_{j}\in class\left\{{l}\right\}$
11:     Compute
$\widetilde{x}_{j+1}=\widetilde{y}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{y}_{j+1}\right)$
12:end for
13:$x=P^{T}\widetilde{x}$.
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times inner product and then orthogonal partitioning largely avoids this problem. On the one hand, the computation time of inner product is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
4 Convergence analysis
Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the preprocessed linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left\|P^{T}\widetilde{x}_{j}-P^{T}\widetilde{x}_{*}\right\|_{2}^{2}=\left\|P^{T}\left(\widetilde{x}_{j}-\widetilde{x}_{*}\right)\right\|_{2}^{2}=\left\|\widetilde{x}_{j}-\widetilde{x}_{*}\right\|_{2}^{2}.$ (4.1)
It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
Lemma 1.19 Assumes that a system of linear equations with a coefficient matrix of $\widetilde{A}\in{{\mathbb{R}}^{m\times m}}$ and a right-hand vector of $\widetilde{f}\in{{\mathbb{R}}^{m}}$ is compatible. Using the idea of uniform partitioning, the row vectors of the coefficient matrix $\widetilde{A}$ are partitioned to obtain $A_{\tau_{1}},A_{\tau_{2}},...,A_{\tau_{k+1}}$. Singular value decomposition is performed on the submatrix $A_{\tau_{i}}\in{{\mathbb{R}}^{m_{\tau_{i}}\times m}}$
$A_{\tau_{i}}=U_{\tau_{i}}\Sigma_{\tau_{i}}V_{\tau_{i}}^{H}$ (4.2)
where $U_{\tau_{i}}$ is a $m_{\tau_{i}}\times m_{\tau_{i}}$ unitary matrix. $\Sigma_{\tau_{i}}$ is a $m_{\tau_{i}}\times m$ diagonal matrix. $V_{\tau_{i}}^{H}$, the conjugate transpose of $V_{\tau_{i}}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{\tau_{i}}$ are singular values of $A_{\tau_{i}}$. Then $A_{\tau_{i}}^{\dagger}=V_{\tau_{i}}\Sigma_{\tau_{i}}^{H}U_{\tau_{i}}^{H}$, and $A_{\tau_{i}}^{\dagger}A_{\tau_{i}}=V_{\tau_{i}}MV_{\tau_{i}}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{\tau_{i}}$.
Lemma 2. Assumes that $V$ is a unitary matrix of $m\times m$ order, which can be considered as a combination of column vectors, denoted as $V=[v_{1},v_{2},...,v_{m}]$ then
(1) $v_{i}^{H}v_{j}=\left\{\begin{aligned} 0,~{}~{}~{}~{}~{}i\neq j\\ 1,~{}~{}~{}~{}~{}i=j\end{aligned}\right.$
(2) For any $m$ dimensional vector $u$, it can be uniquely represented as a column linear combination $u=\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}$, and the Euclidean norm $\left\|u\right\|_{2}^{2}=\left\|\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}\right\|_{2}^{2}=\sum_{i=1}^{m}\alpha_{i}^{2}$.
Similar to Ref. 20 , we have the following convergence theorem.
Theorem 1.For the linear system (1.1) $Ax=f$, the iteration sequence $\left\{x_{j}\right\}_{j=0}^{\infty}$ generated by POBK algorithm GREBK(k) converges to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
Proof. By POBK algorithm GREBK(k), the main calculation is for block wise linear systems in the following form
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right).$
For convenience, denote $A_{\tau_{j1}}=A_{\tau_{0}}$. For the submatrix $A_{\tau_{0}}$ of $\widetilde{A}$, according to Lemma 1, there are $A_{\tau_{0}}=U_{\tau_{0}}\Sigma_{\tau_{i}}V_{\tau_{0}}^{H}$, $A_{\tau_{0}}^{\dagger}=V_{\tau_{0}}\Sigma_{\tau_{i}}^{H}U_{\tau_{0}}^{H}$, and therefore $A_{\tau_{0}}^{\dagger}A_{\tau_{0}}=V_{\tau_{0}}MV_{\tau_{0}}^{H}$, $M=diag\left(1,...,1,0,...,0\right)$, where the number of 1 is the rank of $A_{\tau_{0}}$, then
$x_{1}-x_{*}$ $=x_{0}+A_{\tau_{0}}^{\dagger}\left(f_{\tau_{0}}-A_{\tau_{0}}x_{0}\right)-x_{*}$ (4.3)
$=x_{0}-x_{*}+A_{\tau_{0}}^{\dagger}\left(A_{\tau_{0}}x_{*}-A_{\tau_{0}}x_{0}\right)$
$=\left(I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\right)\left(x_{0}-x_{*}\right).$
Due to $I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}$ being an orthogonal projection operator and $x_{0}-x_{*}$ being linearly represented by the column vectors of $V_{\tau_{0}}$, i.e.,
$x_{0}-x_{*}=\alpha_{\tau_{0}1}v_{\tau_{0}1}+\alpha_{\tau_{0}2}v_{\tau_{0}2}+...+\alpha_{\tau_{0}m}v_{\tau_{0}m}=V_{\tau_{0}}\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}.$ (4.4)
So
$\left\|x_{1}-x_{*}\right\|_{2}^{2}$ $=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$ (4.5)
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|V_{\tau_{0}}MV_{\tau_{0}}^{H}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|M\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|\left[\alpha_{\tau_{0}1},...,\alpha_{\tau_{0}r},0,...,0\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\gamma_{\tau_{0}}\left\|x_{0}-x_{*}\right\|_{2}^{2}$
$=\left(1-\gamma_{\tau_{0}}\right)\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}$
where $\gamma_{\tau_{0}}=\sum_{i=1}^{r}\left(\alpha_{\tau_{0}i}\right)^{2}/\sum_{i=1}^{m}\left(\alpha_{\tau_{0}i}\right)^{2}$, obviously $10<\gamma_{\tau_{0}}<$ 1. Therefore
$\centering\|x_{j}-x_{\star}\|_{2}^{2}=(1-\gamma_{j})\|x_{j-1}-x_{\star}\|_{2}^{2}<\|x_{j-1}-x_{\star}\|_{2}^{2}.\@add@centering$ (4.6)
i.e., $\mathop{\lim}\limits_{j\to\infty}{\left\|{{x_{j}}-{x_{\star}}}\right\|_{2}}=0$. POBK algorithm GREBK(k) converges. ∎
For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes see Figure 4:
Figure 4:  Two cases of the initial value $x_{0}$ on a 2D System.
Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above situations, we have that
$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.7)
thus
$\left\|x_{j}-x_{*}\right\|_{2}=\left\|x_{j-1}-x_{*}\right\|_{2}cos\theta=...=\left\|x_{1}-x_{*}\right\|_{2}cos^{j-1}\theta.$ (4.8)
According to Theorem 1, so
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left(1-\gamma_{\tau_{0}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.9)
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the hyperplane angle $\theta$. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between the two blocks in $class\left\{i\right\}$ is close to 90 degrees, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $class\left\{l\right\}$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate diagonal matrix, the blocks in the matrix have better orthogonality than the original matrix $A$, and the convergence rate is also faster than the GRBK, RBK (k), and GREBK (k) algorithms.
5 Numerical experiments
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK (k), GREBK (k), and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental testing are all selected from 20 sparse matrices with application backgrounds in the $SuiteSparse\;Matrix\;Collection$, and have different scales (most of which are large sparse matrices). Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the system of equations. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector.
For all the above algorithms, we set the initial value $x_{0}=0$ and the exact solution is $x_{*}$. If the relative error $RSE$ under the current iteration is less than $61e-$ 6 or the number of iterations exceeds $55e+$ 5, the calculation will be stopped, where
$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}}{\left\|x_{*}\right\|_{2}}.$
All experiments were conducted in $MATLAB(R2019a)$ on a PC with AMD Ryzen 7 5825U with Radeon Graphics 2.00 GHz. We use iteration steps (IT) and CPU time in seconds (CPU) to evaluate the numerical performance of different algorithms. Here, IT and CPU are the average iteration steps and CPU time of running the underlying algorithm 10 times. If the corresponding algorithm fails to converge under the specified relative error bound or maximum number of iterations, the corresponding IT and CPU are recorded as Inf and NAN in the table.
We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 1. The definition of matrix density is
$density=\frac{number~{}of~{}nonzeros~{}of~{}an~{}m\times m~{}matrix}{m\cdot m}.$
Table 1:    Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
$Name$ $Kind$ $m$ $Non~{}Num$ $Cond$ $density\left(\%\right)$
$Poli$ $Economic~{}Problem$ 4 00840084\,0084 008 8 18881888\,1888 188 3.115e+02 0.051
$Muu$ $Structural~{}Problem$ 7 01270127\,0127 012 170 134170134170\,134170 134 7.654e+01 0.337
$Poli3$ $Economic~{}Problem$ 16 9551695516\,95516 955 37 8493784937\,84937 849 5.253e+02 0.013
$Ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2 87028702\,8702 870 23 7542375423\,75423 754 7.515e+02 0.288
$Kim1$ $2D/3D~{}Problem$ 38 4153841538\,41538 415 933 195933195933\,195933 195 9.863e+03 0.063
$Bodyy4$ $Structural~{}Problem$ 17 5461754617\,54617 546 121 550121550121\,550121 550 8.059e+02 0.039
$Qpband$ $Optimization~{}Problem$ 20 0002000020\,00020 000 45 0004500045\,00045 000 6.436e+00 0.011
$Rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
$Obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$Spmsrtls$ $Statistical/Mathematical~{}Problem$ 29 9952999529\,99529 995 229 947229947229\,947229 947 1.416e+05 0.026
$Blckhole$ $Structural~{}Problem$ 2 13221322\,1322 132 14 8721487214\,87214 872 4.167e+03 0.327
$Linverse$ $Statistical/Mathematical~{}Problem$ 11 9991199911\,99911 999 95 9779597795\,97795 977 3.947e+03 0.067
$Torsion1$ $Duplicate~{}Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$Polilarge$ $Economic~{}Problem$ 15 5751557515\,57515 575 33 0743307433\,07433 074 2.983e+01 0.014
$Jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
$Dixmaanl$ $Optimization~{}Problem$ 60 0006000060\,00060 000 299 998299998299\,998299 998 1.285e+05 0.008
$Bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
$Crystm03$ $Materials~{}Problem$ 24 6962469624\,69624 696 583 770583770583\,770583 770 2.640e+02 0.096
$Chem97ztz$ $Statistical/Mathematical~{}Problem$ 2 54125412\,5412 541 7 36173617\,3617 361 2.472e+02 0.114
$Wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
The numerical experimental results in Table 2 show that our proposed Kaczmarz (POBK) algorithm, which has been preprocessed and reordered into blocks, is much faster than the Kaczmarz (GRBK) and k-means direct block Kaczmarz (GRBK, GREBK (K), RBK (k)) algorithms of the random block algorithm in terms of iteration steps and computation time, and the improvement rate of algorithm iteration times and computation time is mostly above 90%. It is worth noting that for matrices in the complex field, such as the Kim1 matrix, the GRBK, GREBK (K), and RBK (k) algorithms encountered problems during partitioning and were unable to continue running. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
Table 2:   IT and CPU performance of 20 test matrices under different algorithms.
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $POBK$ $Increase~{}Speed\left(\%\right)$
$Poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 $75$ $\left[99.81,99.90\right]$
CPU $40.63$ $43.43$ $28.71$ $0.15$ $\left[99.48,99.65\right]$
$Muu$ IT $606$ $176$ $269$ $1$ $\left[99.43,99.83\right]$
CPU $2.91$ $7.84$ $2.44$ $0.26$ $\left[89.34,96.68\right]$
$Poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 1 03410341\,0341 034 $\left[98.21,98.61\right]$
CPU $136.48$ $338.26$ $125.88$ $8.92$ $\left[92.91,97.36\right]$
$Ex29$ IT $253$ $10$ $116$ $8$ $\left[20.00,96.84\right]$
CPU $0.19$ $1.37$ $0.20$ $0.12$ $\left[58.33,91.24\right]$
$Kim1$ IT $NAN$ $NAN$ $NAN$ $52$ $\left[NAN,NAN\right]$
CPU $NAN$ $NAN$ $NAN$ $18.27$ $\left[NAN,NAN\right]$
$Bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 $296$ $\left[99.75,99.94\right]$
CPU $NAN$ $NAN$ $807.00$ $30.64$ $\left[96.20,98.42\right]$
$Qpband$ IT $627$ $335$ $292$ $1$ $\left[99.66,99.84\right]$
CPU $1.28$ $194.37$ $4.33$ $1.01$ $\left[21.09,99.48\right]$
$Rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 1 42214221\,4221 422 $\left[99.41,99.72\right]$
CPU $NAN$ $3094.14$ $3074.96$ $78.34$ $\left[95.17,97.47\right]$
$Obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $132$ $\left[97.37,99.65\right]$
CPU $158.43$ $1108.05$ $62.96$ $19.08$ $\left[69.70,98.28\right]$
$Spmsrtls$ IT $Inf$ $Inf$ $Inf$ $2$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $2.05$ $\left[99.91,99.98\right]$
$Blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 1 38113811\,3811 381 $\left[93.78,99.62\right]$
CPU $456.49$ $32.55$ $261.30$ $10.66$ $\left[67.25,97.66\right]$
$Linverse$ IT $Inf$ $Inf$ $Inf$ $10$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $0.54$ $\left[99.97,99.98\right]$
$Torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $132$ $\left[88.94,99.65\right]$
CPU $166.10$ $966.06$ $147.52$ $18.54$ $\left[87.43,98.08\right]$
$Polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 763363763363\,763363 763 5 12851285\,1285 128 $\left[98.59,98.97\right]$
CPU $NAN$ $1002.59$ $813.27$ $23.91$ $\left[97.06,98.05\right]$
$Jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 3 60336033\,6033 603 $\left[98.58,99.28\right]$
CPU $NAN$ $272.82$ $533.79$ $16.24$ $\left[94.05,97.78\right]$
$Dixmaanl$ IT 38 8753887538\,87538 875 $980$ 2 55825582\,5582 558 $9$ $\left[99.08,99.98\right]$
CPU $237.74$ $1687.74$ $17.05$ $0.63$ $\left[96.30,99.96\right]$
$Bcsstm39$ IT $20$ $20$ $20$ $1$ $\left[95.00,95.00\right]$
CPU $0.10$ $5189.52$ $19.14$ $0.04$ $\left[60.00,99.99\right]$
$Crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $1$ $\left[99.86,99.98\right]$
CPU $41.89$ $645.50$ $24.83$ $2.24$ $\left[90.98,99.65\right]$
$Chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $9$ $\left[99.75,99.90\right]$
CPU $3.88$ $5.08$ $2.03$ $0.05$ $\left[97.54,99.02\right]$
$Wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $160$ $\left[94.94,96.77\right]$
C PU $20.78$ $2193.32$ $36.76$ $20.57$ $\left[1.01,99.06\right]$
Figure 5:  The number of iteration steps (left) and average computation time (right) for testing matrix $Poli3$ to reach convergence under different iteration methods.
Figure 6:  The number of iteration steps (left) and average computation time (right) for testing matrix $Polilarge$ to reach convergence under different iteration methods.
Figure 7:  The number of iteration steps (left) and average computation time (right) for testing matrix $Jagmesh4$ to reach convergence under different iteration methods.
Figure 8:  The number of iteration steps (left) and average computation time (right) for testing matrix $Rajat07$ to reach convergence under different iteration methods.
Figure 9:  The number of iteration steps (left) and average computation time (right) for testing matrix $Muu$ to reach convergence under different iteration methods.
5.1 The Number of matrix blocks
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. Moreover, due to the differences in the results of each clustering of the k-means algorithm, the Kaczmarz algorithm, which is reflected in numerical experiments as a k-means clustering block, has significant differences in iteration time even for the same k-value, and the stability of the algorithm is not very good. In our proposed method, the selection of the number of blocks $k$ is also involved. Although it also needs to be set in advance, the value of $k$ can be determined based on the number of conditions in the matrix and the preprocessed ”width”. Numerical experiments have shown that the algorithm has strong stability and will not have significant differences in iteration time for the same $k$ value. Taking the $ex29$ matrix as an example, Figure 10 shows the iteration time changes of GRBK, RBK(k), GREBK(k), and POBK algorithms after 10 iterations with $k$ constant. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in computation time, but also performs better in computation time stability.
Figure 10:  Changes in iteration time for 10 iterations of different algorithms under constant $k$.
5.2 The Robustness of POBK Algorithm
Another advantage of our proposed algorithm is its robustness. Compared to GRBK, RBK(k), and GREBK(k) algorithms, even if different $k$ values are taken, the convergence rate of POBK algorithm does not change much, and it is still better than GRBK, RBK(k), and GREBK(k) algorithms. Taking $Poli$, $Muu$, $Crystm03$, $Torsion1$ matrices as examples, we compared the iteration time and steps required for convergence of GRBK, RBK(k), GREBK(k) algorithms with POBK algorithms at different values. The $k$ in POBK(k) represents the number of blocks, as detailed in Table 3 and Figure 11.
Table 3:   The IT and CPU required for the convergence of GRBK, RBK (k), GREBK (k) algorithms and POBK algorithms with different $k$ values.
$Poli$ $Muu$ $Crystm03$ $Torsion1$
$GRBK$ IT 78 9727897278\,97278 972 $716$ 4 $34$\,3434 343 37 4243742437\,42437 424
CPU $43.04$ $3.38$ $44.20$ $166.10$
$RBK(k)$ IT 36 2233622336\,22336 223 $213$ $863$ 1 19411941\,1941 194
CPU $27.10$ $11.56$ $860.76$ $966.06$
$GREBK(k)$ IT 51 6835168351\,68351 683 $410$ 1 03010301\,0301 030 5 04350435\,0435 043
CPU $27.60$ $3.32$ $16.48$ $147.52$
$POBK(3)$ IT $75$ $1$ $1$ $108$
CPU $0.16$ $0.24$ $2.30$ $34.07$
$POBK(5)$ IT $165$ $7$ $9$ $114$
CPU $0.36$ $0.62$ $7.56$ $26.76$
$POBK(10)$ IT $250$ $8$ $11$ $9122$
CPU $0.80$ $0.55$ $8.29$ $24.21$
$POBK(15)$ IT $792$ $8$ $16$ $132$
CPU $3.16$ $0.49$ $9.82$ $18.54$
$POBK(20)$ IT $851$ $9$ $10$ $151$
CPU $4.27$ $0.52$ $5.75$ $20.47$
Figure 11:  Comparison chart of $CPU$ required for convergence of $GRBK$, $RBK(k)$, $GREBK(k)$ algorithms and $POBK$ algorithms with different $k$ values.
6 Conclusions
This article proposes a new preprocessed orthogonal block wise Kaczmarz (POBK) algorithm based on random block Kaczmarz and k-means technology. We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate block diagonal matrix. Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and calculate the inner product between the centroid coordinates to measure the orthogonality between different blocks, so that we can group the orthogonal blocks into one class. We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence. The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), and GREBK(k) algorithms in terms of iteration steps and CPU time. The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution. At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning. We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition matrices, in order to improve the numerical performance of the POBK algorithm. This is a valuable work worth studying in the future.
Acknowledgements. The authors are very much indebted to the referees for their constructive comments and valuable suggestions. The authors are partially supported by National Natural Science Foundation of China (11101071).

### v3

@@
-Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology
+Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology
 Abstract
 In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$.
-Firstly, the Reverse Cuthill McKee Algorithm (RCM) algorithm is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide orthogonal blocks into one category, which greatly reduces computational complexity.
-In addition, the convergence of POBK algorithm is theoretically proven, and a corresponding analysis of its faster convergence is also provided.
-Finally, numerical simulations indicate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k) and GREBK(k) in both iteration steps (IT) and CPU time aspects.
+Firstly, the sparse Reverse Cuthill-McKee Algorithm (RCM) is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide the mutually orthogonal blocks into one category, which greatly reduces computational complexity.
+In addition, we provide a theoretical proof of the convergence of the POBK algorithm and analyze its faster convergence.
+Finally, numerical simulations demonstrate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k), GREBK(k), and aRBK in terms of CPU time.
 keywords:
-Kaczmarz algorithm, Preprocessing techniques, Orthogonal partitioning, Robustness, k-means
+Kaczmarz algorithm, Sparse Reverse Cuthill-McKee ordering, Preprocessing techniques, Orthogonal partitioning
 journal:
 1 Introduction
-This paper mainly considers the solution of a large uniformly sparse linear system in the form of
+This paper mainly considers the solution of a large sparse linear system in the form of
 $Ax=f$ (1.1)
 where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$.
-One of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method.
-The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i}=\{x\mid\left\langle A_{i},x\right\rangle=f_{i}\}$ defined by the system rows 1 ; 2 ; 19 .
-Specifically, the initial value is set to $x_{0}$, the iteration format of the classic Kaczmarz method is
-$x_{j+1}=x_{j}+\frac{f_{i}-A_{i}^{T}x}{\left\|A_{i}\right\|_{2}^{2}}\left(A_{i}^{T}\right).$ (1.2)
+Nowadays, one of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method [1, 2].
+The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i_{k}}=\{x\in{\mathbb{R}^{m}}\mid A_{(i_{k})}x=f_{i_{k}}\}$ defined by the system rows [1, 2, 25].
+Specifically, the initial value is set to $x_{0}$, assuming that the $i_{k}$th row has been selected at the $j$th iteration, then the $(j+1)$th estimate vector $x_{j+1}$ is obtained by:
+$x_{j+1}=x_{j}+\frac{f_{i_{k}}-A_{(i_{k})}x_{j}}{\left\|A_{(i_{k})}\right\|_{2}^{2}}A_{(i_{k})}^{T}.$ (1.2)
 The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory.
-Therefore, Strohmer and Vershynin 3 ; 4 proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate.
+Therefore, Strohmer and Vershynin [3, 4] proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate.
 However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal.
-Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method 5 .
+Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method [5].
 Nutini et al.
-proposed the greedy distance Kaczmarz (GDK) method 6 .
+proposed the greedy distance Kaczmarz (GDK) method [6].
 Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set.
-They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version, which can converge to the system’s solution faster.
+They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version [7, 8], which can converge to the system’s solution faster.
 Liu et al.
-proposed a variant of GRK to solve the ridge regression problem 9 .
+proposed a variant of GRK to solve the ridge regression problem [9].
 Du Yishu et al.
-introduced an effective distance based row selection criterion and proposed the distance based greedy random Kaczmarz (GDRK) algorithm 10 , which has better solving performance than traditional RK and GRK algorithms.
-The block Kaczmarz (BK) method 11 ; 12 is a natural extension of the classical Kaczmarz method and another effective way to improve its convergence performance.
-Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row metrics in each iteration.
-Given an initial solution $x_{0}$, in the k-th iteration, first select a subset of travel indicators $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$.
-Using the pseudo inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
+introduced an effective distance-based row selection criterion and proposed the distance-based greedy random Kaczmarz (GDRK) algorithm [10], which has better solving performance than traditional RK and GRK algorithms.
+The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method and provides another effective way to improve its convergence performance.
+Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration.
+Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$.
+Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
 $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right),$ (1.3)
 where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$.
-Needell et al.
-proposed the Good Saving and Block Random Kaczmarz (RBK) algorithms 13 ; 14 ; 15 .
-Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) 18 .
+Recently, Needell et al.
+proposed the good saving and block random Kaczmarz (RBK) algorithms [13, 14, 15].
+Necoara developed a family of randomized block Kaczmarz algorithms that use a subset of constrained and extrapolated steps at each step, and can be deployed on distributed computing units [16].
+Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) [19].
 Zhang et al.
-proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual 16 .
+proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual [17].
 Based on the k-means clustering algorithm, Jiang, Li et al.
-proposed the row clustering block Kaczmarz (RBK (k)) algorithm based on the k-means algorithm 17 and the global random block Kaczmarz method 18 , further improving the convergence rate of the algorithm.
-On this basis, Zheng Wenhao et al.
-explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK (k)) algorithm 20 , which has a much faster computational speed than the RBK (k) method.
+proposed the row clustering block Kaczmarz (RBK(k)) algorithm based on the k-means algorithm [18] and the global random block Kaczmarz method [19], further improving the convergence rate of the algorithm.
+On this basis, Zheng et al.
+explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK(k)) algorithm [20], which has a much faster computational speed than the RBK(k) method.
+In 2024, Zhang et al.
+combined fast projection and weight averaging techniques to propose the aRBK algorithm [21].
 In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm.
 However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm.
-When faced with sparsely distributed matrices, the iteration time for the block Kaczmarz algorithm to obtain accurate solutions is not ideal.
-To improve the convergence speed of block Kaczmarz and the convergence of iterative sequences, this paper constructs an orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques.
-The algorithm mainly uses preprocessing techniques to rearrange the original matrix, making the new matrix an approximate block diagonal matrix.
-Then, the matrix is divided into blocks, with the aim of grouping orthogonal blocks into one category.
-In this way, each iteration in each category undergoes two orthogonal projections on two mutually orthogonal hyperplanes, so the iterative solution can approach the exact solution more quickly.
+For the sparse matrix with a scattered distribution, the iteration time for obtaining an exact solution using the block Kaczmarz algorithm is not ideal.
+To improve the convergence speed of block Kaczmarz, this paper constructs a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing technology.
+The proposed algorithm utilizes preprocessing techniques to rearrange the original matrix so that it becomes approximately a banded matrix.
+Then, the matrix is divided into several blocks, with the goal of grouping orthogonal blocks into one category.
+In this way, each iteration in this category is performed as two orthogonal projections on nearly orthogonal hyperplanes, so that the iterative solution converges faster to the exact solution.
 This article theoretically proves the convergence of the POBK algorithm.
-The numerical experimental results also show that compared with the GRBK, RBK(k), and GREBK(k) algorithms, the POBK algorithm is more efficient and ideal in terms of iterative sequence convergence and algorithm running time.
-The structure of this article is as follows.
-In Section 2, we briefly introduced the GRBK algorithm, RBK (k) algorithm, and GREBK (k) algorithm.
-In Section 3, we introduced the preprocessing techniques and partitioning ideas used, analyzed the reasons for the selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm.
+The numerical experimental results also show that compared with the GRBK, RBK(k), GREBK(k) and aRBK algorithms, the POBK algorithm is more efficient and ideal in terms of algorithm running time.
+The structure of this article is as follows: In Section 2, we briefly introduced the GRBK, RBK(k), GREBK(k) and aRBK algorithms.
+In Section 3, we introduce the preprocessing techniques and partitioning ideas used, analyzed the reasons for their selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm.
 In Section 4, we proved the convergence of the proposed method.
-In Section 5, we selected appropriate matrices for numerical simulation to demonstrate the effectiveness of our method.
+In Section 5, we selected some matrices for numerical simulation to demonstrate the effectiveness of our method.
 Finally, in Section 6, we provide conclusions and point out some future directions for work.
@@
 The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems.
-The block Kaczmarz method under greedy algorithms and random strategies is a natural development of the block Kaczmarz method.
+The block Kaczmarz method under greedy algorithms and random strategies (GRBK) is a natural development of the block Kaczmarz method.
 The specific description of the algorithm can be found in Algorithm 1.
-Algorithm 1 GRBK Algorithm 17
+Algorithm 1 GRBK Algorithm [18]
 1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
 2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
-3:Randomly block $A$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ for row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$.
+3:Randomly obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$.
 $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
 In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al.
-first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK (k)) method 17 .
+first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK(k)) method [18].
 Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class.
-In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block sub-matrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm.
+In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block submatrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm.
 The specific description of the RBK(k) algorithm can be found in Algorithm 2.
-Algorithm 2 RBK(k) Algorithm 17
+Algorithm 2 RBK(k) Algorithm [18]
 1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
@@
 10:end for
-In 2023, Zheng Wenhao, et al.
-discovered the relationship between ordinary residuals and standardized residuals 20 , that is, the standardized residuals are the 2-norm of the row vector of the ordinary residuals divided by the corresponding sparse matrix $A$
-$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{i}^{T}\right\|_{2}}=\frac{f_{i}-A_{i}^{T}x_{0}}{\left\|A_{i}^{T}\right\|_{2}}.$ (2.1)
+In 2024, Zheng Wenhao, et al.
+discovered the relationship between ordinary residuals and standardized residuals [20], that is, the standardized residuals are the 2-norm of the row vector of the corresponding sparse matrix $A$ divided by the ordinary residuals.
+$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{(i)}\right\|_{2}}=\frac{f_{i}-A_{(i)}x_{0}}{\left\|A_{(i)}\right\|_{2}}.$ (2.1)
 When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
@@
 and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm.
-So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the blocks of $A$.
-They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to block the standardized residuals clustering, as shown in Algorithm 3.
-Algorithm 3 GREBK(k) Algorithm 20
+So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the partitioning strategy for matrix $A$.
+They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to the block standardized residuals clustering, as shown in Algorithm 3.
+Algorithm 3 GREBK(k) Algorithm [20]
 1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
@@
 10:end for
-3 Orthogonal Blocking Kaczmarz Method Based on Preprocessing Technology
-3.1 Preprocessing Technology–Reverse Cuthill McKee Algorithm
-As is well known, the Kaczmarz algorithm has good convergence for linear systems, especially for sparse linear systems.
-However, Kaczmarz algorithm does not exhibit good convergence speed for all sparse matrices, especially for matrices with large dimensions or scattered structural distributions.
-Even using block clustering and greedy algorithms cannot effectively change this situation.
-We analyze the iterative characteristics of the Kaczmarz algorithm and the structure of matrices in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed and the angle between the hyperplane where the iteration point is located and the next projection hyperplane are closely related.
-Obviously, the smaller the angle between the two hyperplanes, the more iterations, and the slower the convergence.
-The closer the angle is to the perpendicular, the fewer iterations, and the faster the convergence.
+In 2024, Zhang et al.
+[21] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques.
+Their combination can balance convergence speed, convergence range, and computational complexity.
+Please refer to Algorithm 4 for more details.
+Algorithm 4 aRBK Algorithm [21]
+1:$A$, $f$, $x_{0}$, $\left\{\omega_{j}\right\}_{j\in\tau}$, $\alpha$, $K$, $k=0$
+2:An estimation of the unique solution $x_{*}$ to $Ax=f$.
+3:The application definition divides the system into $k$ groups $A_{j}$ and $f_{j}$, $j\in\tau$.
+4:while $k<K$ do
+5:     $k=k+1$
+6:     for $j\in\tau$ do, perform independent sampling $S_{j}$ for each subsystem, then update
+7:         $x_{k}^{j}=x_{k-1}+\alpha\left(A_{S_{j}}^{j}\right)^{\dagger}\left(f_{S_{j}}^{j}-A_{S_{j}}^{j}x_{k-1}\right)$
+8:     end for
+9:     Compute $x_{k}=\sum_{j=1}^{\tau}\omega_{j}x_{k}^{j}$
+10:end while
+3 Orthogonal Block Kaczmarz Method Based on Preprocessing Technology
+3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
+The Kaczmarz algorithm is well known for its good convergence in linear systems, especially for sparse matrices.
+However, it does not exhibit fast convergence speed for all sparse matrices, particularly those with large dimensions or scattered structural distributions.
+Even the use of block clustering and greedy algorithms cannot effectively improve this situation.
+We analyze the iterative characteristics of the Kaczmarz algorithm and the matrix structure in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed is closely related to the angle between the current iteration point’s hyperplane and the next projection one.
+It is evident that a smaller angle between these two hyperplanes leads to more iterations and slower convergence, while an angle closer to perpendicular results in fewer iterations and faster convergence.
 Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases.
@@
 The red dots where the hyperplanes intersect represent the exact solution $x_{*}$.
-The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$, It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration.
+The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$.
+It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non-orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration.
 So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other?
-If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see 21 .
+If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see [22].
 Figure 1:  Comparison of Kaczmarz Iterations on Small and Vertical Angles on a 2D System.
-The Reverse Cuthill McKee (RCM) algorithm was initially used for reordering symmetric matrices.
+Based on the above analysis, it is considered to be cost-effective to make the original matrix into a matrix whose rows are orthogonal to each other by preprocessing the linear system.
+Common sparse matrix ordering algorithms include Column Count Reordering, Reverse Cuthill-McKee algorithm (RCM), Minimum Degree Reordering, Nested Dissection and Spectral Reordering, etc.
+[26, 27, 28, 29, 30, 31].
+Each of these algorithms has its own advantages and is suitable for different types of computational tasks.
+Despite the various methods of ordering available, experimental comparisons have shown that the RCM algorithm usually produces superior computational results in terms of reducing bandwidth and profile of a sparse matrix.
+The main objective of the RCM algorithm is to minimize the maximum distance between non-zero elements and the main diagonal, thereby reducing bandwidth and greatly improving the orthogonality between matrix row blocks.
+Consequently, for the block Kaczmarz algorithm, the RCM algorithm is the recommended choice.
+The RCM algorithm [26] was initially used for reordering symmetric matrices.
 It will generate a matrix with rows and columns numbered according to the node numbers.
 By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated.
-The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., $\widetilde{A}$ has diagonal elements closer to the diagonal.
-Fortunately, the RCM algorithm also performs equally well on asymmetric matrices.
-So, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm, transforming the original coefficient starting matrix into an approximate (block) diagonal matrix, see Figure 2.
-Obviously, there are many rows or row blocks that are mutually orthogonal in the approximate (block) diagonal matrix, see Figure 3.
-And this improves the drawback of the original matrix structure being arranged randomly.
-According to our above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
+The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., the nonzero elements of $\widetilde{A}$ are closer to the diagonal and the run-time complexity is proved to be linear in the number of nonzeros in the matrix [26].
+Fortunately, the RCM algorithm also often performs equally well on asymmetric matrices (see http://www.ece.northwestern.edu/local-apps/matlabhelp/techdoc/ref/symrcm.html).
+Therefore, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm to transform the original coefficient matrix into an approximate banded matrix.
+Obviously, in the approximate banded matrix, there are many rows or blocks of rows that are orthogonal to each other, see Figure 2, which improves the disadvantage of the random arrangement of the original matrix structure.
+Based on the above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
+Figure 2:   The orthogonality representation of the blckhole matrix after being rearranged using the RCM algorithm.
 In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/).
@@
 Figure 3:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
-3.2 Proposed methods
+Note that, the run-time complexity of the RCM algorithm is proved to be linear in the number of nonzeros in the matrix [26].
+Therefore, the addition of the RCM algorithm does not lead to significant time overhead.
+We select five matrices muu, poli3, blckhole, kim1, and wathen100 of different dimensions as examples, and count the time consumption of the RCM algorithm for preprocessing them, as shown in Table 1.
+Table 1:    Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
+$Name$ $m$ $density\left(\%\right)$ $RCM\;preprocessing\;time\;consumption(s)$
+$muu$ 7 01270127\,0127 012 $0.337$ 0.002220
+$poli3$ 16 9551695516\,95516 955 $0.013$ 0.002261
+$blckhole$ 2 13221322\,1322 132 $0.327$ 0.000558
+$kim1$ 38 4153841538\,41538 415 $0.063$ 0.018929
+$wathen100$ 30 4013040130\,40130 401 $0.051$ 0.005443
+3.2 The Proposed methods
 As mentioned above, both RBK(k) and GREK(k) algorithms use the k-means algorithm to group matrix rows.
@@
 Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
-Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$, and the preprocessed matrix is
+Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$ as follows.
 $PAP^{T}$ (3.1)
-where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements in the original coefficient matrix $A$ closer to the main diagonal elements.
-Even if the matrix structure is closer to block diagonal form and $P$ is a permutation matrix (i.e., with $P^{T}P=E$), the linear system (1.1) becomes
+where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements of the original coefficient matrix $A$ closer to the main diagonal, i.e., it makes the structure of matrix $A$ closer to the banded matrix form, and $P$ is a substitution matrix (i.e., there is $P^{T}P=E$), so the linear system (1.1) becomes
 $PAP^{T}Px=Pf.$ (3.2)
 Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
-Since $\widetilde{A}$ is an approximate block diagonal matrix, as long as $\widetilde{A}$ is appropriately partitioned, orthogonality between blocks can be achieved, see Figure 3.
-Figure 3:   The orthogonality representation of the matrix Blackhole after being rearranged using the RCM algorithm.
-Consider dividing $\widetilde{A}$ into $k$ blocks, taking $\left[m/k\right]$ elements from each block, and storing the remaining elements in the $k+1$-th block.
-Perform corresponding partitioning on $\widetilde{f}$, that is, partitioning:
-$\widetilde{A}=\begin{bmatrix}A_{\tau_{1}}\\ A_{\tau_{2}}\\ ...\\ A_{\tau_{k}}\\ A_{\tau_{k+1}}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{\tau_{1}}\\ f_{\tau_{2}}\\ ...\\ f_{\tau_{k}}\\ f_{\tau_{k+1}}\\ \end{bmatrix}.$
+Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row chunking of $\widetilde{A}$, see Figure 2.
+Due to the natural orthogonality of banded matrices (see Figure 2), we directly chunk $\widetilde{A}$ uniformly in order.
+Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding chunking is done for $f$,
+$\widetilde{A}=\begin{bmatrix}A_{1}\\ A_{2}\\ ...\\ A_{k}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{1}\\ f_{2}\\ ...\\ f_{k}\\ \end{bmatrix}.$
 As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution.
-So obviously, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non orthogonal hyperplanes.
+Then, according to the analysis in Section 3.1,, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non-orthogonal hyperplanes, see Figure 1.
 So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
-Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{\tau_{i}}$ as $\overline{A}_{\tau_{i}}$, $i=1,2,...,k+1$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates to form an inner product table $G$, $G\left(i,j\right)$ which is used to measure the orthogonality between blocks $\overline{A}_{\tau_{i}}$ and $\overline{A}_{\tau_{j}}$.
-Moreover, as $\widetilde{A}$ is an approximate block diagonal matrix, it is unreasonable and convenient to use the inner product of centroid coordinates to reflect the orthogonality between blocks.
-And we divide the matrix $\widetilde{A}$ into $k+1$ blocks in advance, in fact, we only need to do inner product on the centroid coordinates of these $k+1$ blocks.
-Moreover, since $G$ is a symmetric matrix, we only need to calculate $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times, because $k\ll m$, this is a relatively small number.
-Next, based on our above analysis, our aim is to group two row blocks that are mutually orthogonal, i.e., $G\left(i,j\right)=0$ into one class to accelerate the Kaczmarz method.
-However, it is evident that there may be more than one block orthogonal to any block, which can easily result in some row blocks being in multiple categories simultaneously.
-In order to avoid this situation, we stipulate that for each block $A_{\tau_{i}}$, $i=1,2,..k$, we select the first block in $\left\{A_{\tau_{i}},A_{\tau_{i+1}},...,A_{\tau_{k+1}}\right\}$ that has an inner product of 00 with the centroid coordinates of the current block $A_{\tau_{i}}$ and place it in one category, marking the current class as $class\left\{{i}\right\}$, and ensuring that the blocks between each class are not duplicated.
-For blocks that have not entered any class, in order to avoid non-convergence caused by information loss, we classify all these blocks into a new class, denoted as $class\left\{l\right\}$.
-This regulation is reasonable because $\widetilde{A}$ approximates the properties of block diagonal matrices.
-Obtain solutions through block Kaczmarz iteration method on two categories
+Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $C\left(i,j\right)=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $C\left(i,j\right)$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$.
+And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks.
+And we divide the matrix $\widetilde{A}$ into $k$ blocks beforehand, and we only need to calculate the cosine of the angle between these $k$ blocks, and since $C$ is a symmetric matrix, we only need to calculate $\frac{k\left(k-1\right)}{2}$ times in practice.
+Since $k\ll m$, this is a fairly small number.
+Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $C(i,j)=0$, into one class.
+Here we set a threshold $thr$ such that when $C(i,j)<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal, i.e., so that $C(i,j)=0$.
+However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time.
+In order to avoid this situation, we stipulate that for each block $A_{i}$, $i=1,2,..k$, the first block $A_{j}\in\left\{A_{i},A_{i+1},...,A_{k}\right\}$ orthogonal to $A_{i}$ is selected and placed in a class, noting the current class as $class\left\{i\right\}$, and requiring that there is no duplication of blocks between classes.
+We merge these orthogonal classes, denoted $Oclass$.
+For the blocks that do not go into $Oclass$, in order to avoid the non-convergence situation caused by the loss of information, we assign all of these blocks in a new class, denoted $Nclass$.
+This regulation is reasonable because of the nature of the A-approximate banded matrix.
+The solution is obtained by the block Kaczmarz iterative method on both categories:
 $\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
@@
 Algorithm 5 POBK Algorithm
-1:sparse matrix $A$, right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, the maximum number of iterations $l$.
+1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
 2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
 3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
-4:Apply the k-means method to $\widetilde{A}$ to obtain a partition $\{\tau_{1},...,\tau_{k+1}\}$ of the row $\{1,...m\}$, and obtain $k+1$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k+1$.
-$\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
-5:Calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates of each block to form the inner product table $G$.
-6:According to the above regulations, place mutually orthogonal blocks in $class\left\{i\right\}$, and ensure that the blocks between each class are not duplicated.
-Divide all remaining blocks into a new $class\left\{l\right\}$.
+4:Perform uniform chunking of $\widetilde{A}$ and $\widetilde{f}$ in sequential order, and obtain $k$ groups $A_{i}$ and $f_{i}$, $i=1,2,...,k$.
+$\overline{A}=[\overline{A}_{1},...,\overline{A}_{k}]$, $\overline{f}=[\overline{f}_{1},...,\overline{f}_{k}]$, where $\overline{A}_{i}$ and $\overline{f}_{i}$ respectively represent the center of $A_{i}$ and $f_{i}$.
+5:Compute the cosine of the angle $C\left(i,j\right)$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
+6:According to the above regulations, place all two-by-two mutually orthogonal blocks in a category $Oclass$, place all remaining blocks in a new class $Nclass$.
 7:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
-8:     Select   $\tau_{j1},\tau_{j2}\in class\left\{{i}\right\}$
+8:     Select   $\tau_{1},\tau_{2}\in Oclass$
 9:     Compute
-$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right)$
-$\widetilde{y}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{j2}}^{\dagger}\left(f_{\tau_{j2}}-A_{\tau_{j2}}\widetilde{x}_{{j+1}/2}\right)$
-10:     Select   $\tau_{j}\in class\left\{{l}\right\}$
+$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{1}}^{\dagger}\left(f_{\tau_{1}}-A_{\tau_{1}}\widetilde{x}_{j}\right)$
+$\widetilde{x}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{2}}^{\dagger}\left(f_{\tau_{2}}-A_{\tau 2}\widetilde{x}_{{j+1}/2}\right)$
+10:     Select   $\tau_{j}\in Nclass$
 11:     Compute
-$\widetilde{x}_{j+1}=\widetilde{y}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{y}_{j+1}\right)$
+$\widetilde{x}_{j+2}=\widetilde{x}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j+1}\right)$
 12:end for
@@
 In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal.
-Our proposed idea of performing $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times inner product and then orthogonal partitioning largely avoids this problem.
-On the one hand, the computation time of inner product is very short because $k\ll m$.
+Our proposed idea of performing $\frac{k\left(k-1\right)}{2}$ calculations of the angle cosine values and then orthogonal chunking avoids this problem to a large extent.On the one hand, the computation time of the angle cosine value is very short because $k\ll m$.
 On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
@@
 It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
-Lemma 1.19 Assumes that a system of linear equations with a coefficient matrix of $\widetilde{A}\in{{\mathbb{R}}^{m\times m}}$ and a right-hand vector of $\widetilde{f}\in{{\mathbb{R}}^{m}}$ is compatible.
-Using the idea of uniform partitioning, the row vectors of the coefficient matrix $\widetilde{A}$ are partitioned to obtain $A_{\tau_{1}},A_{\tau_{2}},...,A_{\tau_{k+1}}$.
-Singular value decomposition is performed on the submatrix $A_{\tau_{i}}\in{{\mathbb{R}}^{m_{\tau_{i}}\times m}}$
-$A_{\tau_{i}}=U_{\tau_{i}}\Sigma_{\tau_{i}}V_{\tau_{i}}^{H}$ (4.2)
-where $U_{\tau_{i}}$ is a $m_{\tau_{i}}\times m_{\tau_{i}}$ unitary matrix.
-$\Sigma_{\tau_{i}}$ is a $m_{\tau_{i}}\times m$ diagonal matrix.
-$V_{\tau_{i}}^{H}$, the conjugate transpose of $V_{\tau_{i}}$, is a unitary matrix of $m\times m$ order.
-The elements on the diagonal of matrix $\Sigma_{\tau_{i}}$ are singular values of $A_{\tau_{i}}$.
-Then $A_{\tau_{i}}^{\dagger}=V_{\tau_{i}}\Sigma_{\tau_{i}}^{H}U_{\tau_{i}}^{H}$, and $A_{\tau_{i}}^{\dagger}A_{\tau_{i}}=V_{\tau_{i}}MV_{\tau_{i}}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{\tau_{i}}$.
-Lemma 2. Assumes that $V$ is a unitary matrix of $m\times m$ order, which can be considered as a combination of column vectors, denoted as $V=[v_{1},v_{2},...,v_{m}]$ then
-(1) $v_{i}^{H}v_{j}=\left\{\begin{aligned} 0,~{}~{}~{}~{}~{}i\neq j\\ 1,~{}~{}~{}~{}~{}i=j\end{aligned}\right.$
-(2) For any $m$ dimensional vector $u$, it can be uniquely represented as a column linear combination $u=\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}$, and the Euclidean norm $\left\|u\right\|_{2}^{2}=\left\|\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}\right\|_{2}^{2}=\sum_{i=1}^{m}\alpha_{i}^{2}$.
-Similar to Ref. 20 , we have the following convergence theorem.
-Theorem 1.For the linear system (1.1) $Ax=f$, the iteration sequence $\left\{x_{j}\right\}_{j=0}^{\infty}$ generated by POBK algorithm GREBK(k) converges to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
-Proof. By POBK algorithm GREBK(k), the main calculation is for block wise linear systems in the following form
-$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right).$
-For convenience, denote $A_{\tau_{j1}}=A_{\tau_{0}}$.
-For the submatrix $A_{\tau_{0}}$ of $\widetilde{A}$, according to Lemma 1, there are $A_{\tau_{0}}=U_{\tau_{0}}\Sigma_{\tau_{i}}V_{\tau_{0}}^{H}$, $A_{\tau_{0}}^{\dagger}=V_{\tau_{0}}\Sigma_{\tau_{i}}^{H}U_{\tau_{0}}^{H}$, and therefore $A_{\tau_{0}}^{\dagger}A_{\tau_{0}}=V_{\tau_{0}}MV_{\tau_{0}}^{H}$, $M=diag\left(1,...,1,0,...,0\right)$, where the number of 1 is the rank of $A_{\tau_{0}}$, then
-$x_{1}-x_{*}$ $=x_{0}+A_{\tau_{0}}^{\dagger}\left(f_{\tau_{0}}-A_{\tau_{0}}x_{0}\right)-x_{*}$ (4.3)
-$=x_{0}-x_{*}+A_{\tau_{0}}^{\dagger}\left(A_{\tau_{0}}x_{*}-A_{\tau_{0}}x_{0}\right)$
-$=\left(I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\right)\left(x_{0}-x_{*}\right).$
-Due to $I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}$ being an orthogonal projection operator and $x_{0}-x_{*}$ being linearly represented by the column vectors of $V_{\tau_{0}}$, i.e.,
-$x_{0}-x_{*}=\alpha_{\tau_{0}1}v_{\tau_{0}1}+\alpha_{\tau_{0}2}v_{\tau_{0}2}+...+\alpha_{\tau_{0}m}v_{\tau_{0}m}=V_{\tau_{0}}\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}.$ (4.4)
-So
-$\left\|x_{1}-x_{*}\right\|_{2}^{2}$ $=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$ (4.5)
-$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|V_{\tau_{0}}MV_{\tau_{0}}^{H}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$
-$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|M\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}\right\|_{2}^{2}$
-$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|\left[\alpha_{\tau_{0}1},...,\alpha_{\tau_{0}r},0,...,0\right]^{H}\right\|_{2}^{2}$
-$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\gamma_{\tau_{0}}\left\|x_{0}-x_{*}\right\|_{2}^{2}$
-$=\left(1-\gamma_{\tau_{0}}\right)\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}$
-where $\gamma_{\tau_{0}}=\sum_{i=1}^{r}\left(\alpha_{\tau_{0}i}\right)^{2}/\sum_{i=1}^{m}\left(\alpha_{\tau_{0}i}\right)^{2}$, obviously $10<\gamma_{\tau_{0}}<$ 1. Therefore
-$\centering\|x_{j}-x_{\star}\|_{2}^{2}=(1-\gamma_{j})\|x_{j-1}-x_{\star}\|_{2}^{2}<\|x_{j-1}-x_{\star}\|_{2}^{2}.\@add@centering$ (4.6)
-i.e., $\mathop{\lim}\limits_{j\to\infty}{\left\|{{x_{j}}-{x_{\star}}}\right\|_{2}}=0$. POBK algorithm GREBK(k) converges. ∎
-For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes see Figure 4:
+Lemma 1 [25]. Assumes that the singular value decomposition is performed on the submatrix $A_{i}\in{{\mathbb{R}}^{m_{i}\times m}}$,
+$A_{i}=U_{i}\Sigma_{i}V_{i}^{H}$ (4.2)
+where $U_{i}$ is a $m_{i}\times m_{i}$ unitary matrix.
+$\Sigma_{i}$ is a $m_{i}\times m$ diagonal matrix.
+$V_{i}^{H}$, the conjugate transpose of $V_{i}$, is a unitary matrix of $m\times m$ order.
+The elements on the diagonal of matrix $\Sigma_{i}$ are singular values of $A_{i}$.
+Then the Moore–Penrose pseudoinverse of $A_{i}$ $A_{i}^{\dagger}=V_{i}\Sigma_{i}^{H}U_{i}^{H}$, and $A_{i}^{\dagger}A_{i}=V_{i}MV_{i}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{i}$.
+Lemma 2 [13]. For any $A\in{{\mathbb{R}}^{n\times m}}$, we have
+$\sigma_{min}(A^{\dagger})=\frac{1}{\sigma_{max}(A)},$
+where $\sigma_{min}\left(.\right)$ denotes the smallest nonzero singular value of a matrix and $\sigma_{max}\left(.\right)$ denotes the largest singular value of a matrix.
+Similar to Ref. [16], we have the following convergence theorem.
+Theorem 1.
+If the linear system (1.1) $Ax=f$ is consistent, then the POBK algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
+In particular, we have the following linear convergence rate in expectation:
+$\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{jk}\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.3)
+Proof. When $\tau_{j}\in Nclass$, the POBK algorithm involves one block per iteration and computes the linear system by the following form
+${x}_{j+1}={x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}{x}_{j}\right).$ (4.4)
+According to Lemma 1, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$.
+Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
+$x_{j+1}-x_{*}$ $=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}$ (4.5)
+$=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)$
+$=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).$
+Therefore, we have
+$\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$ (4.6)
+$=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$
+$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.$
+Taking the conditional expectation on both sides of inequality (4.6), since the matrix $\widetilde{A}$ is uniformly partitioned, the probability of each block being selected is equal, i.e., $p_{\tau}=\frac{1}{k}$, we get
+$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]$ $=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$ (4.7)
+$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$
+$\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}$
+$\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},$
+and assuming that $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
+$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.8)
+By taking the full expectation on both sides of the above equation, we have
+$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.9)
+When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the POBK algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
+$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}$ (4.10)
+$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
+We divide the matrix into $k$ blocks, where $k_{1}$ blocks are partitioned into $Oclass$ and $k_{2}$ blocks are partitioned into $Nclass$, and $k_{1}+k_{2}=k$.
+According to our POBK algorithm, iterating first in $Oclass$ and then in $Nclass$, then it is clear that the iteration is done once and
+$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}$ (4.11)
+$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
+The inequality (4.3) follows by induction on $j$. ∎
+Remark 1.
+As can be seen from the above proof, the POBK algorithm converges as expected and the convergence rate is related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks.
+The power of the convergence rate of the POBK algorithm is much larger than the power of the general Kaczmarz algorithm, and theoretically there will be a much faster convergence speed.
+In order to visualize the convergence speed of the POBK algorithm more intuitively, below we explain the reason why the POBK algorithm converges faster from the perspective of the hyperplane pinch angle.
+For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes, see Figure 4.
 Figure 4:  Two cases of the initial value $x_{0}$ on a 2D System.
-Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above situations, we have that
+Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above cases, we have that
 $cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.12)
@@
 According to Theorem 1, so
-$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left(1-\gamma_{\tau_{0}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.9)
-From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the hyperplane angle $\theta$.
+$\left\|x_{j}-x_{*}\right\|_{2}^{2}\leq\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.14)
+From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the value of the cosine of the hyperplane pinch angle.
 The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm.
@@
 Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
-In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between the two blocks in $class\left\{i\right\}$ is close to 90 degrees, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms.
+In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between two blocks that are mutually orthogonal in $Oclass$ is close to $90^{\circ}$, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms.
 Therefore, the iterative solution will approach the exact solution more quickly.
-Although the blocks in $class\left\{l\right\}$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate diagonal matrix, the blocks in the matrix have better orthogonality than the original matrix $A$, and the convergence rate is also faster than the GRBK, RBK (k), and GREBK (k) algorithms.
+Although the blocks in $Nclass$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate banded matrix, the orthogonality between the blocks in the matrix $\widetilde{A}$ is better than the original matrix $A$.
+Meanwhile, two nearly orthogonal blocks in $Oclass$ have a high probability of being non-adjacent, then the matrices in $Nclass$, although not orthogonal to each other, will have a much lower probability of being adjacent to each other, i.e., the number of times to compute the cosine value of the smaller pinch angle will be much lower, and according to Equation (4.14), its convergence speed is also faster than that of the GRBK, RBK (k), GREBK (k) and aRBK algorithms.
 5 Numerical experiments
-In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK (k), GREBK (k), and POBK methods, verifying the effectiveness and efficiency of the algorithm.
-The matrices used for numerical experimental testing are all selected from 20 sparse matrices with application backgrounds in the $SuiteSparse\;Matrix\;Collection$, and have different scales (most of which are large sparse matrices).
-Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the system of equations.
-We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector.
+5.1 Numerical simulation
+In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK(k), GREBK(k), aRBK and POBK methods, verifying the effectiveness and efficiency of the algorithm.
+The matrices used for numerical experimental tests are 20 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density.
+Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system.
+We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector space.
 For all the above algorithms, we set the initial value $x_{0}=0$ and the exact solution is $x_{*}$.
@@
 $rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
+$tols4000$ $ComputationalFluidDynamics~{}Problem$ 4 00040004\,0004 000 8 78487848\,7848 784 2.356e+07 0.055
 $obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
@@
 $jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
-$Dixmaanl$ $Optimization~{}Problem$ 60 0006000060\,00060 000 299 998299998299\,998299 998 1.285e+05 0.008
 $bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
@@
 $wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
-The numerical experimental results in Table 2 show that our proposed Kaczmarz (POBK) algorithm, which has been preprocessed and reordered into blocks, is much faster than the Kaczmarz (GRBK) and k-means direct block Kaczmarz (GRBK, GREBK (K), RBK (k)) algorithms of the random block algorithm in terms of iteration steps and computation time, and the improvement rate of algorithm iteration times and computation time is mostly above 90%.
-It is worth noting that for matrices in the complex field, such as the Kim1 matrix, the GRBK, GREBK (K), and RBK (k) algorithms encountered problems during partitioning and were unable to continue running.
+Since the GRBK, GREBK(k) and RBK(k) algorithms involve only one block per iteration, yet the aRBK and POBK algorithms involve all the blocks of the matrix per iteration, a direct comparison of the number of iterative steps (IT) is not fair.
+As can be seen from Table 3, even with $k$ times the number of iteration steps of the POBK algorithm($k<20$), it is still far superior to the GRBK, GREBK(k) and RBK(k) algorithms.
+As for the computation time, the POBK algorithm is much faster than the randomized chunking algorithms Kaczmarz (GRBK, aRBK) and the k-means direct chunking Kaczmarz (GREBK(k), RBK(k)) algorithms, and we give a comparison of the average running time of the different algorithms for some of the matrices in Figures 5, 6 and 7.
+It is worth noting that for the matrices over complex field, such as kim1 matrices, the GRBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and could not continue to run.
 However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
+From the convergence analysis in Section 4, it can be seen that the convergence speed of the POBK algorithm is related to the maximum and minimum non-zero singular values of the blocks and the angle of the hyperplane.
+We analyzed the matrix structure, and found that other matrices either have smaller ”bandwidth” after preprocessing to accelerate the convergence of the algorithm, or the gap between the largest and smallest non-zero singular values after preprocessing the chunks becomes smaller to accelerate the convergence of the algorithm.
+However, wathen100 is a very special matrix.
+Its original matrix is presented as a banded form, and the ”bandwidth” of the RCM algorithm after preprocessing is slightly larger than the original matrix bandwidth (see Figure 8).
+At the same time, the difference between the maximum and the minimum non-zero singular values after the chunking is also larger (the minimum is 6, and the maximum can reach 40), which will undoubtedly also make the POBK algorithm converge slowly.
+As a result, the wathen100 matrix has a smaller number of iterations, but the iteration time is comparable to that of the GRBK algorithm.
+In addition, we find that there is almost no difference in the running time of the wathen100 matrix between preprocessing with the RCM algorithm and without reordering, which reinforces our idea.
+In this case (sparse matrices with a narrow banding structure), using the original matrix directly may be a better choice.
+This is also in line with our theme that the POBK algorithm presents good convergence for sparse matrices with scattered arrangements.
 Table 3:   IT and CPU performance of 20 test matrices under different algorithms.
-$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $POBK$ $Increase~{}Speed\left(\%\right)$
-$Poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 $75$ $\left[99.81,99.90\right]$
-CPU $40.63$ $43.43$ $28.71$ $0.15$ $\left[99.48,99.65\right]$
-$Muu$ IT $606$ $176$ $269$ $1$ $\left[99.43,99.83\right]$
-CPU $2.91$ $7.84$ $2.44$ $0.26$ $\left[89.34,96.68\right]$
-$Poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 1 03410341\,0341 034 $\left[98.21,98.61\right]$
-CPU $136.48$ $338.26$ $125.88$ $8.92$ $\left[92.91,97.36\right]$
-$Ex29$ IT $253$ $10$ $116$ $8$ $\left[20.00,96.84\right]$
-CPU $0.19$ $1.37$ $0.20$ $0.12$ $\left[58.33,91.24\right]$
-$Kim1$ IT $NAN$ $NAN$ $NAN$ $52$ $\left[NAN,NAN\right]$
-CPU $NAN$ $NAN$ $NAN$ $18.27$ $\left[NAN,NAN\right]$
-$Bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 $296$ $\left[99.75,99.94\right]$
-CPU $NAN$ $NAN$ $807.00$ $30.64$ $\left[96.20,98.42\right]$
-$Qpband$ IT $627$ $335$ $292$ $1$ $\left[99.66,99.84\right]$
-CPU $1.28$ $194.37$ $4.33$ $1.01$ $\left[21.09,99.48\right]$
-$Rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 1 42214221\,4221 422 $\left[99.41,99.72\right]$
-CPU $NAN$ $3094.14$ $3074.96$ $78.34$ $\left[95.17,97.47\right]$
-$Obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $132$ $\left[97.37,99.65\right]$
-CPU $158.43$ $1108.05$ $62.96$ $19.08$ $\left[69.70,98.28\right]$
-$Spmsrtls$ IT $Inf$ $Inf$ $Inf$ $2$ $\left[99.99,99.99\right]$
-CPU $NAN$ $NAN$ $NAN$ $2.05$ $\left[99.91,99.98\right]$
-$Blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 1 38113811\,3811 381 $\left[93.78,99.62\right]$
-CPU $456.49$ $32.55$ $261.30$ $10.66$ $\left[67.25,97.66\right]$
-$Linverse$ IT $Inf$ $Inf$ $Inf$ $10$ $\left[99.99,99.99\right]$
-CPU $NAN$ $NAN$ $NAN$ $0.54$ $\left[99.97,99.98\right]$
-$Torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $132$ $\left[88.94,99.65\right]$
-CPU $166.10$ $966.06$ $147.52$ $18.54$ $\left[87.43,98.08\right]$
-$Polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 763363763363\,763363 763 5 12851285\,1285 128 $\left[98.59,98.97\right]$
-CPU $NAN$ $1002.59$ $813.27$ $23.91$ $\left[97.06,98.05\right]$
-$Jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 3 60336033\,6033 603 $\left[98.58,99.28\right]$
-CPU $NAN$ $272.82$ $533.79$ $16.24$ $\left[94.05,97.78\right]$
-$Dixmaanl$ IT 38 8753887538\,87538 875 $980$ 2 55825582\,5582 558 $9$ $\left[99.08,99.98\right]$
-CPU $237.74$ $1687.74$ $17.05$ $0.63$ $\left[96.30,99.96\right]$
-$Bcsstm39$ IT $20$ $20$ $20$ $1$ $\left[95.00,95.00\right]$
-CPU $0.10$ $5189.52$ $19.14$ $0.04$ $\left[60.00,99.99\right]$
-$Crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $1$ $\left[99.86,99.98\right]$
-CPU $41.89$ $645.50$ $24.83$ $2.24$ $\left[90.98,99.65\right]$
-$Chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $9$ $\left[99.75,99.90\right]$
-CPU $3.88$ $5.08$ $2.03$ $0.05$ $\left[97.54,99.02\right]$
-$Wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $160$ $\left[94.94,96.77\right]$
-C PU $20.78$ $2193.32$ $36.76$ $20.57$ $\left[1.01,99.06\right]$
-Figure 5:  The number of iteration steps (left) and average computation time (right) for testing matrix $Poli3$ to reach convergence under different iteration methods.
-Figure 6:  The number of iteration steps (left) and average computation time (right) for testing matrix $Polilarge$ to reach convergence under different iteration methods.
-Figure 7:  The number of iteration steps (left) and average computation time (right) for testing matrix $Jagmesh4$ to reach convergence under different iteration methods.
-Figure 8:  The number of iteration steps (left) and average computation time (right) for testing matrix $Rajat07$ to reach convergence under different iteration methods.
-Figure 9:  The number of iteration steps (left) and average computation time (right) for testing matrix $Muu$ to reach convergence under different iteration methods.
-5.1 The Number of matrix blocks
+$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $POBK$
+$poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 1 72117211\,7211 721 $28$
+CPU $40.63$ $43.43$ $28.71$ $33.75$ $0.07$
+$muu$ IT $606$ $176$ $269$ $13$ $1$
+CPU $2.91$ $7.84$ $2.44$ $3.51$ $0.22$
+$poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 $899$ 1 03410341\,0341 034
+CPU $136.48$ $338.26$ $125.88$ $121.63$ $7.38$
+$ex29$ IT $253$ $10$ $116$ $7$ $8$
+CPU $0.19$ $1.37$ $0.20$ $0.51$ $0.10$
+$kim1$ IT $NAN$ $NAN$ $NAN$ $124$ $52$
+CPU $NAN$ $NAN$ $NAN$ $299.93$ $18.27$
+$bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 18 3251832518\,32518 325 $112$
+CPU $NAN$ $NAN$ $807.00$ $3574.54$ $12.97$
+$qpband$ IT $627$ $335$ $292$ $11$ $1$
+CPU $1.28$ $194.37$ $4.33$ $3.36$ $1.01$
+$rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 6 89668966\,8966 896 1 08010801\,0801 080
+CPU $NAN$ $3094.14$ $3074.96$ $944.40$ $57.52$
+$tols4000$ IT $491$ $10$ $927$ $11$ $1$
+CPU $0.29$ $4.63$ $0.49$ $0.25$ $0.04$
+$obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $339$ $127$
+CPU $158.43$ $1108.05$ $62.96$ $195.34$ $18.72$
+$spmsrtls$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
+CPU $NAN$ $NAN$ $NAN$ $NAN$ $2.24$
+$blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 2 30823082\,3082 308 1 05510551\,0551 055
+CPU $456.49$ $32.55$ $261.30$ $151.08$ $10.10$
+$linverse$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
+CPU $NAN$ $NAN$ $NAN$ $NAN$ $0.42$
+$torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $386$ $132$
+CPU $166.10$ $966.06$ $147.52$ $209.02$ $18.57$
+$polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 $63 7$ 69776977\,6977 697 5 12851285\,1285 128
+CPU $NAN$ $1002.59$ $813.27$ $975.98$ $20.40$
+$jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 7 10471047\,1047 104 2 87328732\,8732 873
+CPU $NAN$ $272.82$ $533.79$ $292.11$ $10.40$
+$bcsstm39$ IT $20$ $20$ $20$ $4$ $1$
+CPU $0.10$ $5189.52$ $19.14$ $4.09$ $0.04$
+$crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $29$ $1$
+CPU $41.89$ $645.50$ $24.83$ $29.81$ $2.15$
+$chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $20$ $9$
+CPU $3.88$ $5.08$ $2.03$ $0.61$ $0.04$
+$wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $49$ $150$
+C PU $20.78$ $2193.32$ $36.76$ $30.46$ $20.57$
+Figure 5:  Average computational time for test matrices $poli3$ and $polilarge$ to reach convergence for different iteration methods.
+Figure 6:  Average computational time for test matrices $jagmesh4$ and $rajat07$ to reach convergence for different iteration methods.
+Figure 7:  Average computational time for test matrices $muu$ and $chem97ztz$ to reach convergence for different iteration methods.
+Figure 8:   Comparison of matrix wathen100 before and after RCM preprocessing.
+5.2 The Number of matrix blocks $k$
 When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm.
+According to Equation (4.14), the convergence speed of the POBK algorithm is directly related to the cosine of the angle between row blocks.
+When the number of dimensions $m$ increases, $m>2$.
+Here, the pinch angle $\theta$ involves both mutually orthogonal blocks and non-mutually orthogonal blocks.
+When we place mutually orthogonal blocks in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0.
+At this point, it is important to pay attention to the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the POBK algorithm to be critical, which also determines the number of matrix chunks.
+Here we define two variables: the Orthogonality Proportion $zn$ and the Non-orthogonality Severity $nn$ to serve as selection criteria for the number of chunks.
+Definition 1. For the cosine value table $C$, if the total number of elements is $num=k\times k$, the number of zero elements is $n_{1}$, then the Orthogonality Proportion $zn$ is denoted by
+$zn=\frac{n_{1}}{num}=\frac{n_{1}}{k\times k}.$
+Definition 2.
+For the cosine value table $C$, the total number of elements is $num=k\times k$, the number of non-zero elements is $n_{2}$, and the average of these non-zero elements is $nm$, then the Non-orthogonality Severity $nn$ is denoted by
+$nn=\frac{n_{2}\times nm}{num}=\frac{n_{2}\times nm}{k\times k}.$
+It is worth noting that the variable $nn$ is not directly defined by the number of non-zero elements in the matrix $C$ as a percentage of the total number of elements, because we have to consider not only the number of non-zero elements, but also the size of the values of these non-zero elements.
+If these non-zero elements have large values, it means that the angle between some blocks is small, which also makes the iteration slower, which is exactly what we do not want to see.
+We take the ex29 and wathen100 matrices as examples.
+For the ex29 matrix, from Figure 9, $zn$ increases as $k$ increases, and $nn$ is small at $k=2,3,5$, but the value of $zn$ at $k=2,3$ is too small, so it is appropriate to choose $k=5$, and the lowest point of the running time box-and-line diagram is also at $k=5$ from Figure 11.
+For the wathen100 matrix, from Figure 10, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 11.
+Figure 9: $zn$ and $nn$ values of the ex29 matrix for different values of $k$.
+Figure 10: $zn$ and $nn$ values of the muu matrix for different values of $k$.
+Figure 11:  Boxplots of runtime for ex29 and muu matrices for different values of $k$.
+5.3 Stability of the POBK algorithm
 The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$.
-Moreover, due to the differences in the results of each clustering of the k-means algorithm, the Kaczmarz algorithm, which is reflected in numerical experiments as a k-means clustering block, has significant differences in iteration time even for the same k-value, and the stability of the algorithm is not very good.
-In our proposed method, the selection of the number of blocks $k$ is also involved.
-Although it also needs to be set in advance, the value of $k$ can be determined based on the number of conditions in the matrix and the preprocessed ”width”.
-Numerical experiments have shown that the algorithm has strong stability and will not have significant differences in iteration time for the same $k$ value.
-Taking the $ex29$ matrix as an example, Figure 10 shows the iteration time changes of GRBK, RBK(k), GREBK(k), and POBK algorithms after 10 iterations with $k$ constant.
-It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in computation time, but also performs better in computation time stability.
+The difference in the results of each clustering of the k-means algorithm is reflected in numerical experiments as the block Kaczmarz algorithm for k-means clustering, where there is a significant difference in the iteration time even with the same value of $k$, which can lead to the algorithm’s stability is not very good.
+In our proposed method, given a test matrix $A$, which is preprocessed by the RCM algorithm as an approximate banded matrix $\widetilde{A}$, and $\widetilde{A}$ is unique, and since we perform sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and thus the algorithm is stable for the same $k$ value.
+Numerical experiments also show that the algorithm is stable and there is no significant difference in the iteration time for the same value of $k$.
+Taking the ex29 matrix as an example, Figure 12 shows the iteration time changes of GRBK, RBK(k), GREBK(k), aRBK and POBK algorithms after 10 iterations with the same $k$.
+It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of stability.
+The stronger stability of the POBK algorithm can also be seen from the POBK runtime boxplot in Section 5.3.
 Figure 12:  Changes in iteration time for 10 iterations of different algorithms under constant $k$.
-5.2 The Robustness of POBK Algorithm
-Another advantage of our proposed algorithm is its robustness.
-Compared to GRBK, RBK(k), and GREBK(k) algorithms, even if different $k$ values are taken, the convergence rate of POBK algorithm does not change much, and it is still better than GRBK, RBK(k), and GREBK(k) algorithms.
-Taking $Poli$, $Muu$, $Crystm03$, $Torsion1$ matrices as examples, we compared the iteration time and steps required for convergence of GRBK, RBK(k), GREBK(k) algorithms with POBK algorithms at different values.
-The $k$ in POBK(k) represents the number of blocks, as detailed in Table 3 and Figure 11.
-Table 3:   The IT and CPU required for the convergence of GRBK, RBK (k), GREBK (k) algorithms and POBK algorithms with different $k$ values.
-$Poli$ $Muu$ $Crystm03$ $Torsion1$
-$GRBK$ IT 78 9727897278\,97278 972 $716$ 4 $34$\,3434 343 37 4243742437\,42437 424
-CPU $43.04$ $3.38$ $44.20$ $166.10$
-$RBK(k)$ IT 36 2233622336\,22336 223 $213$ $863$ 1 19411941\,1941 194
-CPU $27.10$ $11.56$ $860.76$ $966.06$
-$GREBK(k)$ IT 51 6835168351\,68351 683 $410$ 1 03010301\,0301 030 5 04350435\,0435 043
-CPU $27.60$ $3.32$ $16.48$ $147.52$
-$POBK(3)$ IT $75$ $1$ $1$ $108$
-CPU $0.16$ $0.24$ $2.30$ $34.07$
-$POBK(5)$ IT $165$ $7$ $9$ $114$
-CPU $0.36$ $0.62$ $7.56$ $26.76$
-$POBK(10)$ IT $250$ $8$ $11$ $9122$
-CPU $0.80$ $0.55$ $8.29$ $24.21$
-$POBK(15)$ IT $792$ $8$ $16$ $132$
-CPU $3.16$ $0.49$ $9.82$ $18.54$
-$POBK(20)$ IT $851$ $9$ $10$ $151$
-CPU $4.27$ $0.52$ $5.75$ $20.47$
-Figure 11:  Comparison chart of $CPU$ required for convergence of $GRBK$, $RBK(k)$, $GREBK(k)$ algorithms and $POBK$ algorithms with different $k$ values.
 6 Conclusions
-This article proposes a new preprocessed orthogonal block wise Kaczmarz (POBK) algorithm based on random block Kaczmarz and k-means technology.
-We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate block diagonal matrix.
-Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and calculate the inner product between the centroid coordinates to measure the orthogonality between different blocks, so that we can group the orthogonal blocks into one class.
+In this paper, we propose a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques.
+We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate banded matrix.
+Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and the cosine of the angle between the center of mass coordinates is computed to measure the orthogonality between the different blocks, so that the orthogonal blocks can be grouped together.
 We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence.
-The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), and GREBK(k) algorithms in terms of iteration steps and CPU time.
+The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), GREBK(k), and aRBK algorithms in terms of CPU time.
 The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution.
 At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning.
-We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition matrices, in order to improve the numerical performance of the POBK algorithm.
+We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition coefficient matrices, in order to improve the numerical performance of the POBK algorithm.
 This is a valuable work worth studying in the future.
@@
 The authors are partially supported by National Natural Science Foundation of China (11101071).
+CONFLICT OF INTEREST STATEMENT
+This study does not have any conflicts to disclose.
+DATA AVAILABILITY STATEMENT
+Our computational data comes from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). These data can be freely downloaded and analyzed.

### v4

@@
-Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology
+OBK-RCM: Accelerated Orthogonal Block Kaczmarz Algorithm via RCM Reordering and Dynamic Grouping for Sparse Linear Systems
 Abstract
-In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$.
-Firstly, the sparse Reverse Cuthill-McKee Algorithm (RCM) is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide the mutually orthogonal blocks into one category, which greatly reduces computational complexity.
-In addition, we provide a theoretical proof of the convergence of the POBK algorithm and analyze its faster convergence.
-Finally, numerical simulations demonstrate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k), GREBK(k), and aRBK in terms of CPU time.
+Existing block Kaczmarz methods face challenges in balancing computational efficiency and convergence for large sparse linear systems with scattered nonzero patterns, due to costly partitioning strategies and non-orthogonal projections.
+In this paper, we propose the orthogonal block Kaczmarz (OBK-RCM) algorithm with the Reverse Cuthill-McKee (RCM), which integrates the RCM reordering with a novel orthogonal block partitioning strategy.
+RCM transforms sparse matrices into banded structures to enhance inter-block orthogonality, while dynamic grouping of mutually orthogonal blocks based on angle cosine thresholds reduces iterative complexity.
+In addition, two extended versions (SOBK-RCM and UOBK-RCM) are proposed to deal with non-square systems by constructing extended matrices without sacrificing sparsity.
+This work offers a practical framework for efficient sparse linear algebra solvers.
+Experiments on 33 real-world and synthetic matrices show that OBK-RCM achieves 10-50 times faster CPU time (up to several hundred) and 50–90% fewer iterations than state-of-the-art methods (RBK, RBK(k), GREBK(k), aRBK), especially for scattered sparse structures in most cases.
+Theoretical analysis confirms linear convergence, driven by hyperplane orthogonality.
 keywords:
-Kaczmarz algorithm, Sparse Reverse Cuthill-McKee ordering, Preprocessing techniques, Orthogonal partitioning
+Kaczmarz algorithm; Sparse linear systems; Reverse Cuthill-McKee ordering; Orthogonal partitioning
 journal:
 1 Introduction
-This paper mainly considers the solution of a large sparse linear system in the form of
-$Ax=f$ (1.1)
-where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$.
-Nowadays, one of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method [1, 2].
+This paper focuses on solving large-scale sparse linear systems of the form
+$Ax=f,$ (1.1)
+where $A\in{{\mathbb{R}}^{m\times n}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{n}}$.
+Nowadays, one of the classic and effective iterative projection methods for solving such systems is the Kaczmarz method [1, 2].
 The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i_{k}}=\{x\in{\mathbb{R}^{m}}\mid A_{(i_{k})}x=f_{i_{k}}\}$ defined by the system rows [1, 2, 20].
@@
 $x_{j+1}=x_{j}+\frac{f_{i_{k}}-A_{(i_{k})}x_{j}}{\left\|A_{(i_{k})}\right\|_{2}^{2}}A_{(i_{k})}^{T}.$ (1.2)
-The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory.
-Therefore, Strohmer and Vershynin [3, 4] proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate.
-However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal.
-Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method [5].
-Nutini et al.
-proposed the greedy distance Kaczmarz (GDK) method [6].
-Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set.
-They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version [7, 8], which can converge to the system’s solution faster.
-Liu et al.
-proposed a variant of GRK to solve the ridge regression problem [9].
-Du Yishu et al.
-introduced an effective distance-based row selection criterion and proposed the distance-based greedy random Kaczmarz (GDRK) algorithm [10], which has better solving performance than traditional RK and GRK algorithms.
-The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method and provides another effective way to improve its convergence performance.
+The classic Kaczmarz method suffers from slow convergence due to its dependence on row selection order and lacks rigorous theoretical guarantees.
+To address these limitations, Strohmer and Vershynin [3, 4] introduced the randomized Kaczmarz (RK) algorithm, achieving an exponential convergence rate in expectation.
+However, RK’s random row selection criterion remains suboptimal [4].
+Subsequent studies proposed greedy variants: Ansorge developed the greedy residual Kaczmarz (GEK) [5], Nutini et al.
+formulated the greedy distance Kaczmarz (GDK) [6], and Bai and Wu designed the greedy random Kaczmarz (GRK) with relaxation [7, 8].
+Further extensions include Liu et al.’s GRK adaptation for ridge regression problem [9] and Du et al.’s distance-based GDRK algorithm [10], which outperforms classical RK and GRK in efficiency.
+The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method.
 Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration.
 Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$.
-Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
+Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is:
 $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right),$ (1.3)
 where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$.
-Recently, Needell et al.
-proposed the good saving and block random Kaczmarz (RBK) algorithms [13, 14, 15].
-Necoara developed a family of randomized block Kaczmarz algorithms that use a subset of constrained and extrapolated steps at each step, and can be deployed on distributed computing units [16].
-Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) [19].
+Building on these advancements, Needell et al.
+introduced the randomized block Kaczmarz (RBK) algorithm [13, 14].
+Subsequent work integrated greedy strategies into RBK, yielding the greedy random block Kaczmarz (GRBK) method [17].
 Zhang et al.
-proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual [17].
-Based on the k-means clustering algorithm, Jiang, Li et al.
-proposed the row clustering block Kaczmarz (RBK(k)) algorithm based on the k-means algorithm [18] and the global random block Kaczmarz method [19], further improving the convergence rate of the algorithm.
-On this basis, Zheng et al.
-explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK(k)) algorithm [20], which has a much faster computational speed than the RBK(k) method.
-In 2024, Zhang et al.
-combined fast projection and weight averaging techniques to propose the aRBK algorithm [21].
-In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm.
-However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm.
-For the sparse matrix with a scattered distribution, the iteration time for obtaining an exact solution using the block Kaczmarz algorithm is not ideal.
-To improve the convergence speed of block Kaczmarz, this paper constructs a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing technology.
-The proposed algorithm utilizes preprocessing techniques to rearrange the original matrix so that it becomes approximately a banded matrix.
-Then, the matrix is divided into several blocks, with the goal of grouping orthogonal blocks into one category.
-In this way, each iteration in this category is performed as two orthogonal projections on nearly orthogonal hyperplanes, so that the iterative solution converges faster to the exact solution.
-This article theoretically proves the convergence of the POBK algorithm.
-The numerical experimental results also show that compared with the GRBK, RBK(k), GREBK(k) and aRBK algorithms, the POBK algorithm is more efficient and ideal in terms of algorithm running time.
-The structure of this article is as follows: In Section 2, we briefly introduced the GRBK, RBK(k), GREBK(k) and aRBK algorithms.
-In Section 3, we introduce the preprocessing techniques and partitioning ideas used, analyzed the reasons for their selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm.
-In Section 4, we proved the convergence of the proposed method.
-In Section 5, we selected some matrices for numerical simulation to demonstrate the effectiveness of our method.
-Finally, in Section 6, we provide conclusions and point out some future directions for work.
+further enhanced residual-based selection with the greedy random Motzkin-Kaczmarz (GRMK) algorithm [15], while Jiang and Li incorporated K-means clustering to develop the row-clustered RBK(k) [16] and global randomized block variants [17].
+Zheng et al.
+refined this framework by analyzing residual standardization, proposing the greedy residual block Kaczmarz (GREBK(k)) [18].
+Most recently, Zhang et al.
+(2024) combined fast projection and weighted averaging in the aRBK algorithm [19], achieving a balance between speed and robustness.
+While block Kaczmarz methods (e.g., RBK [13], GREBK(k) [18], aRBK [19]) improve convergence over row-wise iterations, two critical limitations hinder their efficiency:
+  • 1.
+    High partitioning cost: Existing strategies (e.g., K-means clustering [16]) require $O(mk)$ computations for block division, becoming prohibitive for large-scale sparse matrices.
+  • 2.
+    Slow convergence due to non-orthogonal projections: Random or greedy block selection often results in adjacent blocks with small angles between hyperplanes (Figure 1), leading to redundant iterations.
+To address these gaps, we propose the orthogonal block Kaczmarz (OBK-RCM) algorithm with RCM, featuring three core innovations:
+  • 1.
+    RCM reordering: The Reverse Cuthill-McKee reordering concentrates nonzeros near the diagonal, forming an approximate banded matrix with enhanced inter-block orthogonality (Section 3.1).
+  • 2.
+    Dynamic orthogonal partitioning: Blocks with near-zero angle cosines are grouped into orthogonal classes (Oclass), enabling rapid convergence via sequential orthogonal projections (Section 3.3).
+  • 3.
+    Non-square system compatibility: OBK-RCM extends to handle non-square systems without sacrificing sparsity, through extended matrix construction (SOBK-RCM for $m>n$, UOBK-RCM for $m<n$) (see, Eq.
+4.1–4.3), avoiding information loss (Section 4.0).
+This work bridges the gap between low-cost partitioning and accelerated convergence, offering a unified framework for sparse linear systems.
 2 Introduction to existing block methods
 The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems.
-The block Kaczmarz method under greedy algorithms and random strategies (GRBK) is a natural development of the block Kaczmarz method.
-The specific description of the algorithm can be found in Algorithm 1.
-Algorithm 1 GRBK Algorithm [18]
-1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
-2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
-3:Randomly obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$.
-$\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
-4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
-5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
-6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
-7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
-8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
-9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
-10:end for
-In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al.
-first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK(k)) method [18].
-Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class.
-In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block submatrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm.
-The specific description of the RBK(k) algorithm can be found in Algorithm 2.
 Algorithm 1 RBK(k) Algorithm [15]
@@
 10:end for
-In 2024, Zheng Wenhao, et al.
-discovered the relationship between ordinary residuals and standardized residuals [20], that is, the standardized residuals are the 2-norm of the row vector of the corresponding sparse matrix $A$ divided by the ordinary residuals.
-$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{(i)}\right\|_{2}}=\frac{f_{i}-A_{(i)}x_{0}}{\left\|A_{(i)}\right\|_{2}}.$ (2.1)
-When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
-$d_{i}=\lvert{\hat{r}_{i}}\rvert$ (2.2)
-and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm.
-So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the partitioning strategy for matrix $A$.
-They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to the block standardized residuals clustering, as shown in Algorithm 3.
+Firstly, in 2022, Jiang et al. proposed a randomized block Kaczmarz (RBK(k)) method by combining the K-mean clustering algorithm with greedy randomization technique [15], see Algorithm 1 for details.
+Secondly, in 2024, based on the relationship between common residuals and standardized residuals, Zheng et al.
+[17] proposed the block Kaczmarz algorithm (GREBK(k)) for clustering and chunking of standardized residuals combined with the K-means algorithm, see Algorithm 2 for details.
 Algorithm 2 GREBK(k) Algorithm [17]
@@
 2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
-3:Apply the k-means method to $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$.
+3:Apply the K-means method to the normalized residual vector $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$.
 $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
 10:end for
-In 2024, Zhang et al.
+In the same year, Zhang et al.
 [19] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques.
 Their combination can balance convergence speed, convergence range, and computational complexity.
-Please refer to Algorithm 4 for more details.
+See Algorithm 3 for details.
 Algorithm 3 aRBK Algorithm [19]
@@
 10:end while
-3 Orthogonal Block Kaczmarz Method Based on Preprocessing Technology
-3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
-The Kaczmarz algorithm is well known for its good convergence in linear systems, especially for sparse matrices.
-However, it does not exhibit fast convergence speed for all sparse matrices, particularly those with large dimensions or scattered structural distributions.
-Even the use of block clustering and greedy algorithms cannot effectively improve this situation.
-We analyze the iterative characteristics of the Kaczmarz algorithm and the matrix structure in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed is closely related to the angle between the current iteration point’s hyperplane and the next projection one.
-It is evident that a smaller angle between these two hyperplanes leads to more iterations and slower convergence, while an angle closer to perpendicular results in fewer iterations and faster convergence.
-Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases.
-$H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively.
+These algorithms (detailed in Algorithms 1-3) demonstrate how hybrid strategies - clustering of residuals, standardized error metrics, and projection balancing - overcome traditional Kaczmarz limitations in sparse system solving.
+3 The orthogonal block Kaczmarz (OBK-RCM) algorithm with RCM
+3.1 Matrix reordering to enhance orthogonality
+The Kaczmarz algorithm exhibits good convergence properties for linear systems, particularly with sparse matrices.
+However, its convergence rate slows significantly for large-scale sparse matrices or those with scattered structural distributions, even when enhanced by block clustering and greedy algorithms.
+This limitation arises because the algorithm’s convergence speed depends critically on the angle between consecutive projection hyperplanes during iterations.
+Specifically, smaller angles between hyperplanes necessitate more iterations and slower convergence, whereas near-orthogonal projections (angles close to 90°) yield faster convergence with fewer steps.
+To illustrate, consider a two-dimensional linear system (Figure 1).
+Here, $H_{1}$, $H_{2}$, $H_{3}$ denote the system’s hyperplanes, $x_{0}$ the initial point, and $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ the iterative solutions.
 The red dots where the hyperplanes intersect represent the exact solution $x_{*}$.
-The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$.
-It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non-orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration.
-So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other?
-If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see [22].
+The left panel shows slow convergence when $H_{1}$ and $H_{2}$ form a small angle, while the right panel demonstrates rapid convergence when $H_{1}$ and $H_{3}$ are orthogonal—requiring only two iterations to reach the exact solution $x_{*}$, compared to four for non-orthogonal cases.
+This motivates the exploration of methods to extract or construct mutually orthogonal rows in the system matrix, which could drastically reduce iteration time (see [22]).
 Figure 1:  Comparison of Kaczmarz Iterations on Small and Vertical Angles on a 2D System.
-Based on the above analysis, it is considered to be cost-effective to make the original matrix into a matrix whose rows are orthogonal to each other by preprocessing the linear system.
-Common sparse matrix ordering algorithms include Column Count Reordering, Reverse Cuthill-McKee algorithm (RCM), Minimum Degree Reordering, Nested Dissection and Spectral Reordering, etc.
-[26, 27, 28, 29, 30, 31].
-Each of these algorithms has its own advantages and is suitable for different types of computational tasks.
-Despite the various methods of ordering available, experimental comparisons have shown that the RCM algorithm usually produces superior computational results in terms of reducing bandwidth and profile of a sparse matrix.
-The main objective of the RCM algorithm is to minimize the maximum distance between non-zero elements and the main diagonal, thereby reducing bandwidth and greatly improving the orthogonality between matrix row blocks.
-Consequently, for the block Kaczmarz algorithm, the RCM algorithm is the recommended choice.
-The RCM algorithm [26] was initially used for reordering symmetric matrices.
-It will generate a matrix with rows and columns numbered according to the node numbers.
-By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated.
-The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., the nonzero elements of $\widetilde{A}$ are closer to the diagonal and the run-time complexity is proved to be linear in the number of nonzeros in the matrix [26].
-Fortunately, the RCM algorithm also often performs equally well on asymmetric matrices (see http://www.ece.northwestern.edu/local-apps/matlabhelp/techdoc/ref/symrcm.html).
-Therefore, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm to transform the original coefficient matrix into an approximate banded matrix.
-Obviously, in the approximate banded matrix, there are many rows or blocks of rows that are orthogonal to each other, see Figure 2, which improves the disadvantage of the random arrangement of the original matrix structure.
-Based on the above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
-Figure 2:   The orthogonality representation of the blckhole matrix after being rearranged using the RCM algorithm.
-In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/).
-Taking matrices poli3, blackhole, dixmaanl, and jagmesh4 as examples, we visualized the sparse structure of these matrices and the matrices that were reordered by the RCM algorithm.
-We compared the structural changes before and after processing with the RCM algorithm, as shown in Figure 3.
-The blue part represents the distribution of non-zero elements in the matrix.
-From Figure 3, it is evident that after being rearranged by the RCM algorithm, the non-zero elements of the matrix are closer to the diagonal.
-Therefore, choosing the RCM algorithm for preprocessing is reasonable and effective.
-Figure 3:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
-Note that, the run-time complexity of the RCM algorithm is proved to be linear in the number of nonzeros in the matrix [26].
-Therefore, the addition of the RCM algorithm does not lead to significant time overhead.
-We select five matrices muu, poli3, blckhole, kim1, and wathen100 of different dimensions as examples, and count the time consumption of the RCM algorithm for preprocessing them, as shown in Table 1.
-Table 1:    Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
-$Name$ $m$ $density\left(\%\right)$ $RCM\;preprocessing\;time\;consumption(s)$
-$muu$ 7 01270127\,0127 012 $0.337$ 0.002220
-$poli3$ 16 9551695516\,95516 955 $0.013$ 0.002261
-$blckhole$ 2 13221322\,1322 132 $0.327$ 0.000558
-$kim1$ 38 4153841538\,41538 415 $0.063$ 0.018929
-$wathen100$ 30 4013040130\,40130 401 $0.051$ 0.005443
+It is well known that the row blocks of a banded matrix have good orthogonality between them, and the smaller the bandwidth, the stronger the orthogonality.
+Based on the above analysis, we propose to process the original matrix into an approximate banded matrix by reordering.
+Sparse matrix reordering algorithms, such as Reverse Cuthill-McKee (RCM), Column Counting (Colperm), Minimum Degree (Amd), Nested Dissection (Dissect)[23, 24, 25, 26], are matrix reduction bandwidth efficient processing techniques.
+Although there are several sorting methods available, experiments have shown that the RCM algorithm tends to provide better computational results in terms of bandwidth reduction, improved inter-block orthogonality, and algorithmic time-consumption.
+Therefore, for the block Kaczmarz algorithm, the RCM algorithm is preferred.
+Section 3.2 describes the RCM algorithm and its advantages in detail.
+3.2 Reverse Cuthill-McKee Algorithm
+The inverse Cuthill-McKee algorithm [23] is a reordering method for bandwidth minimization of sparse matrices.
+The algorithm drastically reduces the bandwidth of the matrix and enhances the orthogonality of the blocks by rearranging the order of the columns of the matrix rows and clustering the non-zero elements of the original matrix around the diagonal.
+The RCM, Colperm, Amd, and Dissect algorithms have varying results in reducing bandwidth and enhancing orthogonality.
+We use the matlab built-in function $sprandn$ to randomly generate sparse matrices of different dimensions.
+Table 1 shows the original bandwidths ($bw$) of these sparse matrices, as well as the bandwidths after reordering by the RCM, Colperm, Amd, and Dissect algorithms.
+It can be seen that the RCM algorithm can reduce the matrix bandwidth more effectively than the other three sorting algorithms.
+Table 1:   Comparison of processing matrix bandwidth of different sorting algorithms.
+$Name$ $bw$ $Colperm$ $Amd$ $Dissect$ $RCM$
+$1000\times 1000$ 967 984 970 994 28
+$5000\times 5000$ 4959 4904 4720 4837 32
+$10000\times 10000$ 9779 9813 9769 9514 63
+$15000\times 15000$ 14924 14869 14888 14968 68
+In order to have a clearer view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/).
+Taking matrices $poli3$, $blckhole$, $dixmaanl$, and $jagmesh4$ as examples, and we visualize the original structure of these matrices and the structure of these matrices after the reordering by the RCM algorithm, see Figure 3, where the blue part is the distribution of non-zero elements of the matrices.
+It can be clearly seen that after the RCM algorithm reordering, the nonzero elements of the matrices are closer to the diagonal, and the bandwidth of the matrices is significantly reduced.
+Figure 2:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM Reordering.
+Meanwhile, the RCM algorithm is simple and efficient with low time complexity.
+That is, adding the RCM algorithm does not introduce significant time overhead.
+Taking the matrices $muu$, $poli3$, $rajat07$, $blckhole$, and $jagmesh4$ as an example, the time-consuming statistics of the RCM algorithm for reordering them are shown in the Table 2, where $bw$ denotes the original bandwidth of the matrices, and $bw1$ denotes the bandwidth of the matrices after reordering them by RCM.
+It can be seen that RCM has obvious advantages in reducing bandwidth and time consumption.
+For details on the time consumption of the RCM algorithm for reordering sparse matrices, see Ref.
+[23].
+Table 2:    Bandwidth variation and time consumption of the RCM algorithm for reordering sparse matrices.
+$Name$ $m$ $density\left(\%\right)$ $bw$ $bw1$ $time~{}consumption(s)$
+$muu$ 7102 0.337 4696 311 0.0020
+$poli3$ 16955 0.013 16917 5487 0.0022
+$rajat07$ 14842 0.029 1483 413 0.0011
+$blckhole$ 2132 0.327 1805 105 0.0006
+$jagmesh4$ 1440 0.458 1408 56 0.0008
+The above results show that the RCM sorting algorithm not only significantly reduces the bandwidth of the matrix, but also has a very low time overhead.
+Therefore, it is the preferred solution to directly use the RCM sorting algorithm to optimize the structure of sparse linear systems.
 3.3 The Proposed methods
@@
 However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high.
-Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
-Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$ as follows.
-$PAP^{T}$ (3.1)
-where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements of the original coefficient matrix $A$ closer to the main diagonal, i.e., it makes the structure of matrix $A$ closer to the banded matrix form, and $P$ is a substitution matrix (i.e., there is $P^{T}P=E$), so the linear system (1.1) becomes
+Therefore, in order to solve these problems, this paper proposes a simple and fast block Kaczmarz algorithm that combines matrix reordering techniques and orthogonal partitioning ideas.
+Based on the above analysis, We reorder the original matrix $A$ as follows:
+$PAP^{T},$ (3.1)
+where $P$ is obtained by the RCM reordering algorithm and $P$ is a substitution matrix (i.e., $P^{T}P=E$), so the linear system (1.1) becomes
 $PAP^{T}Px=Pf.$ (3.2)
 Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
-Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row chunking of $\widetilde{A}$, see Figure 2.
-Due to the natural orthogonality of banded matrices (see Figure 2), we directly chunk $\widetilde{A}$ uniformly in order.
-Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding chunking is done for $f$,
+Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row blocking of $\widetilde{A}$, see Figure 3.
+Figure 3:   The orthogonality representation of the blckhole matrix after reordering with the RCM algorithm.
+Due to the natural orthogonality of banded matrices, we directly chunk $\widetilde{A}$ uniformly in order.
+Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding blocking is done for $\widetilde{f}$,
 $\widetilde{A}=\begin{bmatrix}A_{1}\\ A_{2}\\ ...\\ A_{k}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{1}\\ f_{2}\\ ...\\ f_{k}\\ \end{bmatrix}.$
-As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution.
-Then, according to the analysis in Section 3.1,, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non-orthogonal hyperplanes, see Figure 1.
-So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
-Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $C\left(i,j\right)=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $C\left(i,j\right)$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$.
+As is well known, the Kaczmarz method iteratively projects the current estimate orthogonally onto successive hyperplanes to approach the exact solution.
+As analyzed in Section 3.1, orthogonal hyperplanes enable accelerated convergence: pairwise orthogonal projections (e.g., $H_{1}\bot H_{3}$ in Figure 1) achieve quadratic convergence within two iterations, while non-orthogonal configurations require significantly more steps.
+This motivates a critical innovation: grouping mutually orthogonal hyperplanes into block projections to eliminate redundant iterations and accelerate convergence—a strategy central to our proposed method.
+Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $cos\theta_{ij}=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $cos\theta_{ij}$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$.
 And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks.
@@
 Since $k\ll m$, this is a fairly small number.
-Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $C(i,j)=0$, into one class.
-Here we set a threshold $thr$ such that when $C(i,j)<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal, i.e., so that $C(i,j)=0$.
+Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $cos\theta_{ij}\approx 0$, into one class.
+Here we set a threshold $thr$ such that when $cos\theta_{ij}<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal.
 However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time.
@@
 This regulation is reasonable because of the nature of the A-approximate banded matrix.
-The solution is obtained by the block Kaczmarz iterative method on both categories:
-$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
+The solution is obtained by the block Kaczmarz method on both categories:
+$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right),$ (3.3)
 performing two iterations on pairwise orthogonal blocks greatly accelerates the convergence speed of the Kaczmarz algorithm, and then performing Kaczmarz iterations on non-orthogonal blocks ensures that matrix information is not omitted.
-Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$.
-Algorithm 5 POBK Algorithm
-1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
+Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$. See Algorithm 4 for details.
+Algorithm 4 OBK-RCM Algorithm
+1:$A$, $f$, $x_{0}$, $k$, $thr$, $l$.
 2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
-3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
+3:Reordering matrices using the RCM algorithm, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
 4:Perform uniform chunking of $\widetilde{A}$ and $\widetilde{f}$ in sequential order, and obtain $k$ groups $A_{i}$ and $f_{i}$, $i=1,2,...,k$.
 $\overline{A}=[\overline{A}_{1},...,\overline{A}_{k}]$, $\overline{f}=[\overline{f}_{1},...,\overline{f}_{k}]$, where $\overline{A}_{i}$ and $\overline{f}_{i}$ respectively represent the center of $A_{i}$ and $f_{i}$.
-5:Compute the cosine of the angle $C\left(i,j\right)$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
-6:According to the above regulations, place all two-by-two mutually orthogonal blocks in a category $Oclass$, place all remaining blocks in a new class $Nclass$.
+5:Compute the cosine of the angle $cos\theta_{ij}$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
+6:Chunking $Oclass$ and $Nclass$ according to the algorithm description.
 7:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
@@
 13:$x=P^{T}\widetilde{x}$.
-Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals.
+Unlike the RBK, RBK(k) and GREBK(k) methods, the OBK-RCM algorithm does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals.
 In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal.
@@
 On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
-4 Convergence analysis
-Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the preprocessed linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
+3.4 Convergence analysis
+Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the reordered linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
 $\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left\|P^{T}\widetilde{x}_{j}-P^{T}\widetilde{x}_{*}\right\|_{2}^{2}=\left\|P^{T}\left(\widetilde{x}_{j}-\widetilde{x}_{*}\right)\right\|_{2}^{2}=\left\|\widetilde{x}_{j}-\widetilde{x}_{*}\right\|_{2}^{2}.$ (3.4)
-It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
-Lemma 1 [25]. Assumes that the singular value decomposition is performed on the submatrix $A_{i}\in{{\mathbb{R}}^{m_{i}\times m}}$,
-$A_{i}=U_{i}\Sigma_{i}V_{i}^{H}$ (4.2)
-where $U_{i}$ is a $m_{i}\times m_{i}$ unitary matrix.
-$\Sigma_{i}$ is a $m_{i}\times m$ diagonal matrix.
-$V_{i}^{H}$, the conjugate transpose of $V_{i}$, is a unitary matrix of $m\times m$ order.
-The elements on the diagonal of matrix $\Sigma_{i}$ are singular values of $A_{i}$.
-Then the Moore–Penrose pseudoinverse of $A_{i}$ $A_{i}^{\dagger}=V_{i}\Sigma_{i}^{H}U_{i}^{H}$, and $A_{i}^{\dagger}A_{i}=V_{i}MV_{i}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{i}$.
-Lemma 2 [13]. For any $A\in{{\mathbb{R}}^{n\times m}}$, we have
-$\sigma_{min}(A^{\dagger})=\frac{1}{\sigma_{max}(A)},$
-where $\sigma_{min}\left(.\right)$ denotes the smallest nonzero singular value of a matrix and $\sigma_{max}\left(.\right)$ denotes the largest singular value of a matrix.
-Similar to Ref. [16], we have the following convergence theorem.
+It can be seen that the convergence analysis of the original linear system and the reordered linear system is consistent, therefore, we will not differentiate this in the future.
 Theorem 1.
-If the linear system (1.1) $Ax=f$ is consistent, then the POBK algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
+If the linear system (1.1) $Ax=f$ is consistent, then the OBK-RCM algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
 In particular, we have the following linear convergence rate in expectation:
 $\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{jk}\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (3.5)
-Proof. When $\tau_{j}\in Nclass$, the POBK algorithm involves one block per iteration and computes the linear system by the following form
+Proof. When $\tau_{j}\in Nclass$, the OBK-RCM algorithm involves one block per iteration and computes the linear system by the following form
 ${x}_{j+1}={x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}{x}_{j}\right).$ (3.6)
-According to Lemma 1, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$.
+According to the singular value decomposition, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$.
 Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
-$x_{j+1}-x_{*}$ $=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}$ (4.5)
-$=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)$
-$=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).$
+$\begin{split}x_{j+1}-x_{*}&=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}\\ &=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)\\ &=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).\end{split}$ (3.7)
 Therefore, we have
-$\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$ (4.6)
-$=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$
-$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.$
+$\begin{split}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)\\ &=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)\\ &=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.\end{split}$ (3.8)
 Taking the conditional expectation on both sides of inequality (3.8), since the matrix $\widetilde{A}$ is uniformly partitioned, the probability of each block being selected is equal, i.e., $p_{\tau}=\frac{1}{k}$, we get
-$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]$ $=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$ (4.7)
-$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$
-$\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}$
-$\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},$
-and assuming that $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
+$\begin{split}\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]&=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]\\ &=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]\\ &\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},\end{split}$ (3.9)
+where $\sigma_{min}(.)$ denotes the minimum non-zero singular value of a matrix, $\sigma_{max}(.)$ denotes the maximum singular value of a matrix.
+It may be useful to set $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$.
+Then we have
 $\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (3.10)
@@
 $\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (3.11)
-When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the POBK algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
-$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}$ (4.10)
-$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
+When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the OBK-RCM algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
+$\begin{split}\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.\end{split}$ (3.12)
 We divide the matrix into $k$ blocks, where $k_{1}$ blocks are partitioned into $Oclass$ and $k_{2}$ blocks are partitioned into $Nclass$, and $k_{1}+k_{2}=k$.
 According to our POBK algorithm, iterating first in $Oclass$ and then in $Nclass$, then it is clear that the iteration is done once and
-$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}$ (4.11)
-$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
+$\begin{split}\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.\end{split}$ (3.13)
 The inequality (3.5) follows by induction on $j$. ∎
 Remark 1.
-As can be seen from the above proof, the POBK algorithm converges as expected and the convergence rate is related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks.
-The power of the convergence rate of the POBK algorithm is much larger than the power of the general Kaczmarz algorithm, and theoretically there will be a much faster convergence speed.
-In order to visualize the convergence speed of the POBK algorithm more intuitively, below we explain the reason why the POBK algorithm converges faster from the perspective of the hyperplane pinch angle.
+From the above proof, it can be seen that the OBK-RCM algorithm converges as expected, and the rate of convergence is closely related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks.
+When the submatrix orthogonality is enhanced (i.e., $\lambda_{max}^{\tau}\approx\lambda_{min}^{\tau}$), the ratio tends to 1 and the rate of convergence increases significantly.
+This explains the key role of orthogonal chunking of the matrix after reordering in accelerating the convergence.
+According to the analysis in Section 3.1, the influence of the angle between the matrix row blocks on the convergence speed is crucial.
+In order to visualize the convergence speed of the OBK-RCM algorithm more intuitively, the following section explains the reason why the OBK-RCM algorithm converges faster from the perspective of the angle between hyperplanes.
 For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes, see Figure 4.
@@
 Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above cases, we have that
-$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.12)
+$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}},$ (3.14)
 thus
@@
 $\left\|x_{j}-x_{*}\right\|_{2}^{2}\leq\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (3.16)
-From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the value of the cosine of the hyperplane pinch angle.
-The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm.
-On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm.
-Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
-In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between two blocks that are mutually orthogonal in $Oclass$ is close to $90^{\circ}$, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms.
-Therefore, the iterative solution will approach the exact solution more quickly.
-Although the blocks in $Nclass$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate banded matrix, the orthogonality between the blocks in the matrix $\widetilde{A}$ is better than the original matrix $A$.
-Meanwhile, two nearly orthogonal blocks in $Oclass$ have a high probability of being non-adjacent, then the matrices in $Nclass$, although not orthogonal to each other, will have a much lower probability of being adjacent to each other, i.e., the number of times to compute the cosine value of the smaller pinch angle will be much lower, and according to Equation (4.14), its convergence speed is also faster than that of the GRBK, RBK (k), GREBK (k) and aRBK algorithms.
-5 Numerical experiments
-5.1 Numerical simulation
-In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK(k), GREBK(k), aRBK and POBK methods, verifying the effectiveness and efficiency of the algorithm.
-The matrices used for numerical experimental tests are 20 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density.
+From Equation (3.16), it can be seen that the convergence speed of the Kaczmarz algorithm is closely related to the cosine of the hyperplane angle.
+The larger the hyperplane angle $\theta$ is (the closer to $90^{\circ}$), and the closer $cos\theta$ is to 0, the faster the algorithm converges; on the other hand, the smaller the hyperplane angle $\theta$ is (the closer to $0^{\circ}$), and the closer $cos\theta$ is to 1, the slower the convergence speed of the algorithm is.
+The OBK-RCM algorithm utilizes the RCM sorting algorithm to reorder the matrices so that the hyperplanes $H_{1}$ and $H_{2}$ are approximately orthogonal ($cos\theta\approx 0$), and the error is rapidly decayed so that the convergence speed is rapidly improved.
+Briefly, the OBK-RCM algorithm achieves superior convergence over traditional block Kaczmarz methods (RBK, RBK(k), GREBK(k), aRBK) through two synergistic mechanisms:
+  1. 1.
+    RCM-Driven Orthogonality Enhancement: By reordering the matrix into a banded structure (Table 1), OBK-RCM forces inter-block angles toward orthogonality, enabling exact solution convergence in two iterations for orthogonal hyperplanes (Figure 1), versus polynomial-time scaling for non-orthogonal cases.
+  2. 2.
+    Intelligent Block Scheduling: The $Oclass$/$Nclass$ partitioning (Section 3.3) systematically minimizes adjacency of non-orthogonal blocks in $Nclass$.
+Although $Nclass$ blocks lack strict orthogonality, their residual angular cosines remain significantly lower than in the original matrix $A$ (Equation 3.16), as RCM reordering inherently suppresses small-angle configurations (Figure 2).
+This dual strategy–orthogonal acceleration in $Oclass$ and mitigated error propagation in $Nclass$–collectively ensures OBK-RCM’s 10-50 times speedup over baseline methods, as quantified in Table 4.
+3.5 Numerical experiments
+In this section, we conducted some experiments to compare the convergence speed of RBK, RBK(k), GREBK(k), aRBK and OBK-RCM methods, verifying the effectiveness and efficiency of the algorithm.
+The matrices used for numerical experimental tests are 15 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density.
 Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system.
@@
 If the relative error $RSE$ under the current iteration is less than $61e-$ 6 or the number of iterations exceeds $55e+$ 5, the calculation will be stopped, where
-$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}}{\left\|x_{*}\right\|_{2}}.$
+$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}^{2}}{\left\|x_{*}\right\|_{2}^{2}}.$
 All experiments were conducted in $MATLAB(R2019a)$ on a PC with AMD Ryzen 7 5825U with Radeon Graphics 2.00 GHz.
@@
 We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 3.
-The definition of matrix density is
+The definition of matrix density is:
 $density=\frac{number~{}of~{}nonzeros~{}of~{}an~{}m\times m~{}matrix}{m\cdot m}.$
 Table 3:  Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
-$Name$ $Kind$ $m$ $Non~{}Num$ $Cond$ $density\left(\%\right)$
-$poli$ $Economic~{}Problem$ 4 00840084\,0084 008 8 18881888\,1888 188 3.115e+02 0.051
-$muu$ $Structural~{}Problem$ 7 01270127\,0127 012 170 134170134170\,134170 134 7.654e+01 0.337
-$poli3$ $Economic~{}Problem$ 16 9551695516\,95516 955 37 8493784937\,84937 849 5.253e+02 0.013
-$ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2 87028702\,8702 870 23 7542375423\,75423 754 7.515e+02 0.288
-$kim1$ $2D/3D~{}Problem$ 38 4153841538\,41538 415 933 195933195933\,195933 195 9.863e+03 0.063
-$bodyy4$ $Structural~{}Problem$ 17 5461754617\,54617 546 121 550121550121\,550121 550 8.059e+02 0.039
-$qpband$ $Optimization~{}Problem$ 20 0002000020\,00020 000 45 0004500045\,00045 000 6.436e+00 0.011
-$rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
-$tols4000$ $ComputationalFluidDynamics~{}Problem$ 4 00040004\,0004 000 8 78487848\,7848 784 2.356e+07 0.055
-$obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
-$spmsrtls$ $Statistical/Mathematical~{}Problem$ 29 9952999529\,99529 995 229 947229947229\,947229 947 1.416e+05 0.026
-$blckhole$ $Structural~{}Problem$ 2 13221322\,1322 132 14 8721487214\,87214 872 4.167e+03 0.327
-$linverse$ $Statistical/Mathematical~{}Problem$ 11 9991199911\,99911 999 95 9779597795\,97795 977 3.947e+03 0.067
-$torsion1$ $Duplicate~{}Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
-$polilarge$ $Economic~{}Problem$ 15 5751557515\,57515 575 33 0743307433\,07433 074 2.983e+01 0.014
-$jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
-$bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
-$crystm03$ $Materials~{}Problem$ 24 6962469624\,69624 696 583 770583770583\,770583 770 2.640e+02 0.096
-$chem97ztz$ $Statistical/Mathematical~{}Problem$ 2 54125412\,5412 541 7 36173617\,3617 361 2.472e+02 0.114
-$wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
-Since the GRBK, GREBK(k) and RBK(k) algorithms involve only one block per iteration, yet the aRBK and POBK algorithms involve all the blocks of the matrix per iteration, a direct comparison of the number of iterative steps (IT) is not fair.
-As can be seen from Table 3, even with $k$ times the number of iteration steps of the POBK algorithm($k<20$), it is still far superior to the GRBK, GREBK(k) and RBK(k) algorithms.
-As for the computation time, the POBK algorithm is much faster than the randomized chunking algorithms Kaczmarz (GRBK, aRBK) and the k-means direct chunking Kaczmarz (GREBK(k), RBK(k)) algorithms, and we give a comparison of the average running time of the different algorithms for some of the matrices in Figures 5, 6 and 7.
-It is worth noting that for the matrices over complex field, such as kim1 matrices, the GRBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and could not continue to run.
-However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
-From the convergence analysis in Section 4, it can be seen that the convergence speed of the POBK algorithm is related to the maximum and minimum non-zero singular values of the blocks and the angle of the hyperplane.
-We analyzed the matrix structure, and found that other matrices either have smaller ”bandwidth” after preprocessing to accelerate the convergence of the algorithm, or the gap between the largest and smallest non-zero singular values after preprocessing the chunks becomes smaller to accelerate the convergence of the algorithm.
-However, wathen100 is a very special matrix.
-Its original matrix is presented as a banded form, and the ”bandwidth” of the RCM algorithm after preprocessing is slightly larger than the original matrix bandwidth (see Figure 8).
-At the same time, the difference between the maximum and the minimum non-zero singular values after the chunking is also larger (the minimum is 6, and the maximum can reach 40), which will undoubtedly also make the POBK algorithm converge slowly.
-As a result, the wathen100 matrix has a smaller number of iterations, but the iteration time is comparable to that of the GRBK algorithm.
-In addition, we find that there is almost no difference in the running time of the wathen100 matrix between preprocessing with the RCM algorithm and without reordering, which reinforces our idea.
-In this case (sparse matrices with a narrow banding structure), using the original matrix directly may be a better choice.
-This is also in line with our theme that the POBK algorithm presents good convergence for sparse matrices with scattered arrangements.
-Table 3:   IT and CPU performance of 20 test matrices under different algorithms.
-$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $POBK$
-$poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 1 72117211\,7211 721 $28$
+$Name$ $Kind$ $m$ $Nonzeros$ $Cond$ $density\left(\%\right)$
+$poli$ $Economic~{}Problem$ 4008 8188 3.115e+02 0.051
+$muu$ $Structural~{}Problem$ 7012 170134 7.654e+01 0.337
+$poli3$ $Economic~{}Problem$ 16955 37849 5.253e+02 0.013
+$ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2870 23754 7.515e+02 0.288
+$kim1$ $2D/3D~{}Problem$ 38415 933195 9.863e+03 0.063
+$qpband$ $Optimization~{}Problem$ 20000 45000 6.436e+00 0.011
+$rajat07$ $Circuit~{}Simulation~{}Problem$ 14842 63913 7.894e+02 0.029
+$blckhole$ $Structural~{}Problem$ 2132 14872 4.167e+03 0.327
+$linverse$ $Statistical/Mathematical~{}Problem$ 11999 95977 3.947e+03 0.067
+$torsion1$ $Duplicate~{}Optimization~{}Problem$ 40000 197608 4.099e+01 0.012
+$polilarge$ $Economic~{}Problem$ 15575 33074 2.983e+01 0.014
+$jagmesh4$ $2D/3D~{}Problem$ 1440 9504 1.516e+04 0.458
+$bcsstm39$ $Structural~{}Problem$ 46772 46772 8.271e+03 0.002
+$crystm03$ $Materials~{}Problem$ 24696 583770 2.640e+02 0.096
+$chem97ztz$ $Statistical/Mathematical~{}Problem$ 2541 7361 2.472e+02 0.114
+Table 4 gives the number of iterations and computation time for these test matrices under different algorithms.
+Since the RBK, GREBK(k) and RBK(k) algorithms involve only one block of data per iteration, whereas the aRBK and OBK-RCM algorithms involve all the blocks of data of the matrix per iteration, it is not fair to directly compare the number of iteration steps (IT) of the different algorithms.
+As can be seen from table 4, even if the OBK-RCM algorithm multiplies the number of iteration steps by $k$ times ($k<20$), it still outperforms the RBK, GREBK(K), and RBK(k) algorithms.
+In terms of computation time, the OBK-RCM algorithm is much faster than the randomized block algorithm Kaczmarz (RBK, aRBK) and the K-means direct block algorithm Kaczmarz(GREBK(k), RBK(k)).
+It is worth noting that for the matrix kim1 in the complex domain, the RBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and therefore could not continue to run.
+However, the OBK-RCM algorithm runs smoothly and achieves the desired convergence rate.
+This shows that it is not only applicable to the real domain but also to the complex domain.
+Table 4:   IT and CPU performance of 15 test matrices under different algorithms
+$Name$ $RBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $OBK-RCM$
+$poli$ IT 74566 38636 51259 1721 28
 CPU 40.63 43.43 28.71 33.75 0.07
-$muu$ IT $606$ $176$ $269$ $13$ $1$
-CPU $2.91$ $7.84$ $2.44$ $3.51$ $0.22$
-$poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 $899$ 1 03410341\,0341 034
-CPU $136.48$ $338.26$ $125.88$ $121.63$ $7.38$
-$ex29$ IT $253$ $10$ $116$ $7$ $8$
-CPU $0.19$ $1.37$ $0.20$ $0.51$ $0.10$
-$kim1$ IT $NAN$ $NAN$ $NAN$ $124$ $52$
-CPU $NAN$ $NAN$ $NAN$ $299.93$ $18.27$
-$bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 18 3251832518\,32518 325 $112$
-CPU $NAN$ $NAN$ $807.00$ $3574.54$ $12.97$
-$qpband$ IT $627$ $335$ $292$ $11$ $1$
-CPU $1.28$ $194.37$ $4.33$ $3.36$ $1.01$
-$rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 6 89668966\,8966 896 1 08010801\,0801 080
-CPU $NAN$ $3094.14$ $3074.96$ $944.40$ $57.52$
-$tols4000$ IT $491$ $10$ $927$ $11$ $1$
-CPU $0.29$ $4.63$ $0.49$ $0.25$ $0.04$
-$obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $339$ $127$
-CPU $158.43$ $1108.05$ $62.96$ $195.34$ $18.72$
-$spmsrtls$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
-CPU $NAN$ $NAN$ $NAN$ $NAN$ $2.24$
-$blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 2 30823082\,3082 308 1 05510551\,0551 055
-CPU $456.49$ $32.55$ $261.30$ $151.08$ $10.10$
+$muu$ IT 160 53 269 13 1
+CPU 8.44 5.38 2.44 3.51 0.22
+$poli3$ IT 26621 6615 18902 1140 1034
+CPU 104.62 125.98 48.07 103.04 7.38
+$ex29$ IT 123 10 116 7 8
+CPU 0.65 1.37 0.20 0.51 0.10
+$kim1$ IT NAN NAN NAN 118 52
+CPU NAN NAN NAN 276.98 18.27
+$qpband$ IT 1687 43 292 11 1
+CPU 3.31 12.33 4.33 3.36 1.01
+$rajat07$ IT Inf 38150 42957 3478 1080
+CPU NAN 2266.82 1889.06 585.08 57.52
+$blckhole$ IT 48327 22213 198250 2308 1055
+CPU 194.91 32.55 188.74 151.08 10.10
 $linverse$ IT Inf Inf Inf Inf 2
 CPU NAN NAN NAN NAN 0.42
-$torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $386$ $132$
-CPU $166.10$ $966.06$ $147.52$ $209.02$ $18.57$
-$polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 $63 7$ 69776977\,6977 697 5 12851285\,1285 128
-CPU $NAN$ $1002.59$ $813.27$ $975.98$ $20.40$
-$jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 7 10471047\,1047 104 2 87328732\,8732 873
-CPU $NAN$ $272.82$ $533.79$ $292.11$ $10.40$
-$bcsstm39$ IT $20$ $20$ $20$ $4$ $1$
-CPU $0.10$ $5189.52$ $19.14$ $4.09$ $0.04$
-$crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $29$ $1$
-CPU $41.89$ $645.50$ $24.83$ $29.81$ $2.15$
-$chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $20$ $9$
-CPU $3.88$ $5.08$ $2.03$ $0.61$ $0.04$
-$wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $49$ $150$
-C PU $20.78$ $2193.32$ $36.76$ $30.46$ $20.57$
-Figure 5:  Average computational time for test matrices $poli3$ and $polilarge$ to reach convergence for different iteration methods.
-Figure 6:  Average computational time for test matrices $jagmesh4$ and $rajat07$ to reach convergence for different iteration methods.
-Figure 7:  Average computational time for test matrices $muu$ and $chem97ztz$ to reach convergence for different iteration methods.
-Figure 8:   Comparison of matrix wathen100 before and after RCM preprocessing.
-5.2 The Number of matrix blocks $k$
-When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm.
-According to Equation (4.14), the convergence speed of the POBK algorithm is directly related to the cosine of the angle between row blocks.
-When the number of dimensions $m$ increases, $m>2$.
-Here, the pinch angle $\theta$ involves both mutually orthogonal blocks and non-mutually orthogonal blocks.
-When we place mutually orthogonal blocks in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0.
-At this point, it is important to pay attention to the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the POBK algorithm to be critical, which also determines the number of matrix chunks.
-Here we define two variables: the Orthogonality Proportion $zn$ and the Non-orthogonality Severity $nn$ to serve as selection criteria for the number of chunks.
+$torsion1$ IT 37424 533 845 386 132
+CPU 166.10 119.36 32.65 209.02 18.57
+$polilarge$ IT 125195 109525 70322 11102 5128
+CPU 411.08 257.46 149.94 625.90 20.40
+$jagmesh4$ IT 385621 125814 73429 7104 2873
+CPU 1173.66 130.63 96.18 292.11 10.40
+$bcsstm39$ IT 50 2 10 4 1
+CPU 0.16 122.52 9.30 4.09 0.04
+$crystm03$ IT 4503 21 196 29 1
+CPU 41.89 36.45 16.20 29.81 2.15
+$chem97ztz$ IT 8727 2037 3574 20 9
+CPU 3.88 3.64 2.03 0.61 0.04
+From the Table 4 and Figure 5, it can be seen that the OBK-RCM algorithm is able to complete the computation within a smaller number of iterations and a shorter CPU time, which exhibits extremely high computational efficiency and stability.
+This indicates that the OBK-RCM algorithm has significant advantages in dealing with large-scale linear systems with sparse square matrices, especially when the structure of the sparse matrices is more decentralized.
+Theoretically, the OBK-RCM algorithm has a better convergence rate by placing $x=P^{T}\widetilde{x}$ inside the loop, which implies that the OBK-RCM algorithm still has potential.
+Figure 5:  Average computational time for test matrices $poli$ and $torsion1$ to reach convergence for different iteration methods.
+Experimental results (Table 4, Figures 5) demonstrate OBK-RCM’s dual superiority:
+  • 1.
+    Accelerated Convergence: Requires 50–90% fewer iterations than RBK/GREBK(k) across matrices with dispersed nonzero patterns (e.g., blckhole: 1,055 vs 198,250 iterations).
+  • 2.
+    Runtime Efficiency: Achieves 10–50 times CPU time (up to several hundred) reduction by eliminating costly K-means clustering and residual sampling (Section 3.3).
+This performance stems from RCM’s banded structure optimization and Oclass/Nclass dynamic partitioning—strategies that jointly maximize hyperplane orthogonality while minimizing small-angle configurations, as quantified by Equation (3.16).
+3.6 Select the optimal number of blocks of the matrix
+As mentioned earlier, for sparse linear systems, blocking is one of the effective means to improve the convergence of kaczmarz algorithm.
+Taking the $crystm03$ matrix as an example, Figure 7 shows IT and CPU required for the matrix to reach convergence for different values of $k$.
+It can be seen that when $k=3$, the OBK-RCM algorithm has the optimal number of iterations and running time.
+Figure 6:  IT and CPU required for matrix $crystm03$ to reach convergence for different values of $k$.
+If $k$ is too large, the dimension of each block becomes smaller, and although the projection computation cost is reduced, the orthogonality between rows of blocks is reduced (i.e., the angle between blocks increases), which slows down convergence; if $k$ is too small, the dimension of a single block is too large, the computational complexity within the block rises and the accelerating effect of the hyperplane orthogonality on the speed of convergence cannot be fully utilized.
+A proper $k$ maximizes the minimum singular value and thus increases the convergence speed.
+According to Equation (3.16), the convergence speed of the OBK-RCM algorithm is directly related to the angle between rows of blocks, where the angle $\theta$ involves both mutually orthogonal and non-mutually orthogonal blocks.
+When mutually orthogonal blocks are grouped in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0.
+At this point, it is important to focus on the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the OBK-RCM algorithm to be crucial to the number of matrix blocks.
+Here, two variables are defined: the orthogonality ratio $zn$ and the degree of non-orthogonality $nn$ to be used as selection criteria for the number $k$ of chunks.
 Definition 1. For the cosine value table $C$, if the total number of elements is $num=k\times k$, the number of zero elements is $n_{1}$, then the Orthogonality Proportion $zn$ is denoted by
@@
 $nn=\frac{n_{2}\times nm}{num}=\frac{n_{2}\times nm}{k\times k}.$
+Theoretically, the optimal $k$ should simultaneously maximize $zn$ and minimize $nn$.
 It is worth noting that the variable $nn$ is not directly defined by the number of non-zero elements in the matrix $C$ as a percentage of the total number of elements, because we have to consider not only the number of non-zero elements, but also the size of the values of these non-zero elements.
 If these non-zero elements have large values, it means that the angle between some blocks is small, which also makes the iteration slower, which is exactly what we do not want to see.
-We take the ex29 and wathen100 matrices as examples.
-For the ex29 matrix, from Figure 9, $zn$ increases as $k$ increases, and $nn$ is small at $k=2,3,5$, but the value of $zn$ at $k=2,3$ is too small, so it is appropriate to choose $k=5$, and the lowest point of the running time box-and-line diagram is also at $k=5$ from Figure 11.
-For the wathen100 matrix, from Figure 10, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 11.
-Figure 9: $zn$ and $nn$ values of the ex29 matrix for different values of $k$.
+We take the $muu$ matrices as examples.
+For the $muu$ matrix, from Figure 7, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 8.
 Figure 7: $zn$ and $nn$ values of the $muu$ matrix for different values of $k$.
-Figure 11:  Boxplots of runtime for ex29 and muu matrices for different values of $k$.
-5.3 Stability of the POBK algorithm
-The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$.
-The difference in the results of each clustering of the k-means algorithm is reflected in numerical experiments as the block Kaczmarz algorithm for k-means clustering, where there is a significant difference in the iteration time even with the same value of $k$, which can lead to the algorithm’s stability is not very good.
-In our proposed method, given a test matrix $A$, which is preprocessed by the RCM algorithm as an approximate banded matrix $\widetilde{A}$, and $\widetilde{A}$ is unique, and since we perform sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and thus the algorithm is stable for the same $k$ value.
-Numerical experiments also show that the algorithm is stable and there is no significant difference in the iteration time for the same value of $k$.
-Taking the ex29 matrix as an example, Figure 12 shows the iteration time changes of GRBK, RBK(k), GREBK(k), aRBK and POBK algorithms after 10 iterations with the same $k$.
-It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of stability.
-The stronger stability of the POBK algorithm can also be seen from the POBK runtime boxplot in Section 5.3.
+Figure 8:  Boxplots of runtime for $muu$ matrices for different values of $k$.
+3.7 Determination of threshold $thr$
+According to the previous definition, the threshold $thr$ is a key parameter in determining the orthogonality of matrix rows and blocks.
+Too large or too small a threshold value will reduce the convergence speed of the OBK-RCM algorithm.
+Therefore, a reasonable choice of the threshold $thr$ requires a trade-off between convergence speed and computational efficiency.
+In this experiment, in order to balance the computational time and computational effectiveness, the running times of typical matrices ($ex29$, $blckhole$, $jagmesh4$ and $linverse$) with different $thr$ values are compared, and the $thr$ value with optimal performance for most of the matrices is selected as the global threshold setting.
+Setting $thr1$, $thr2$, $thr3$, $thr4$, $thr5$, and $thr6$ as $0.5$, $0.2$, $0.1$, $0.05$, $0.02$, and $0.01$, respectively, it can be seen from Table 5 that setting $thr5$ (i.e., setting $thr=0.02$) as the global threshold is worthwhile.
+Table 5:   Comparison of processing matrix bandwidth of different sorting algorithms.
+$Name$ $thr1$ $thr2$ $thr3$ $thr4$ $thr5$ $thr6$
+$ex29$ 0.1126 0.1097 0.1087 0.1086 0.1012 0.1014
+$linverse$ 0.4295 0.4315 0.4281 0.4259 0.4265 0.4281
+$blckhole$ 10.4855 10.257 10.1727 10.2027 10.1444 10.2011
+$jagmesh4$ 10.5939 10.6184 10.5471 10.5426 6810.5386 10.5355
+3.8 Stability of the OBK-RCM algorithm
+The K-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and high efficiency, but the results of this algorithm vary each time the clustering is performed, which is reflected in the numerical experiments as the block Kaczmarz algorithm for K-means clustering, where the computation time is very different even for the same $k$ value, which leads to the algorithm’s stability is not very good.
 Figure 9:  Changes in iteration time for 10 iterations of different algorithms under constant $k$.
-6 Conclusions
-In this paper, we propose a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques.
-We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate banded matrix.
-Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and the cosine of the angle between the center of mass coordinates is computed to measure the orthogonality between the different blocks, so that the orthogonal blocks can be grouped together.
-We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence.
-The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), GREBK(k), and aRBK algorithms in terms of CPU time.
-The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution.
-At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning.
-We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition coefficient matrices, in order to improve the numerical performance of the POBK algorithm.
-This is a valuable work worth studying in the future.
-Acknowledgements.
-The authors are very much indebted to the referees for their constructive comments and valuable suggestions.
-The authors are partially supported by National Natural Science Foundation of China (11101071).
-CONFLICT OF INTEREST STATEMENT
-This study does not have any conflicts to disclose.
+In the method proposed in this section (OBK-RCM), given a test matrix $A$, the approximate banded matrix $A$ after reordering by the RCM algorithm is unique, and since the OBK-RCM algorithm performs sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and hence the algorithm is stable for the same value of $k$.
+Numerical experiments also show that the algorithm is stable, and taking the ex29 matrix as an example, Figure 9 shows the iteration time variations of the RBK, RBK(k), GREBK(k), aRBK, and OBK-RCM algorithms after 10 iterations with $k$ constant.
+It can be clearly seen that the OBK-RCM algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of computation time stability.
+4 Extended Orthogonal Block Kaczmarz for Non-Square Systems
+The OBK-RCM algorithm, introduced in Section 3, achieves efficient solutions for sparse linear systems with scattered nonzero patterns via reordering and orthogonal block partitioning.
+However, a key limitation remains: OBK-RCM requires the coefficient matrix $A$ to be square $\left(i.e.,m=n\right)$.
+To address this, we propose extended matrix formulations that preserve sparsity while extending compatibility to rectangular systems.
+These extensions retain RCM’s orthogonality-enhancing properties and $Oclass/Nclass$ scheduling, ensuring no computational efficiency loss (Figures 10-12).
+Next, the compatible sparse non-square linear system $Ax=f$ is considered, where $A$ is an $m\times n$ matrix $\left(m\neq n\right)$. This section will analyze and discuss both $m>n$ and $m<n$.
+4.1 The $m>n$ case
+For the linear system $Ax=f$, when $m>n$, OBK-RCM fails because the input needs to be square. For this reason, a new matrix $\widehat{A}$ is constructed as:
+$\widehat{A}=[A,O]$ (4.1)
+where $O$ is the $m\times(m-n)$ zero matrix. Thus $\widehat{A}$ becomes an $m\times m$ square matrix.
+Also, let $\widehat{f}=f$, for the equation $\widehat{A}\widehat{x}=\widehat{f}$:
+$\widehat{A}\widehat{x}=\left[A,O\right]\cdot\begin{bmatrix}\widehat{x}_{1}\\ \widehat{x}_{2}\\ \end{bmatrix}=A\widehat{x}_{1}+0^{\prime}=f=\widehat{f}$ (4.2)
+where $\widehat{x}_{1}$ is an $n\times 1$ vector, $\widehat{x}_{2}$ is an $(m-n)\times 1$ vector, $0^{\prime}$ is an $m\times 1$ zero vector.
+Since the linear system $Ax=f$ is compatible, it follows that $\widehat{A}\widehat{x}=\widehat{f}$ is also compatible.
+In fact, the solution of $\widehat{A}\widehat{x}=\widehat{f}$ has a larger solution set than $Ax=f$ because $\widehat{x}_{2}$ can take on any value.
+Based on the above analysis, it can be shown that the first $n$ rows of the solution $\widehat{A}\widehat{x}=\widehat{f}$ is the solution of the linear system $Ax=f$.
+The specific algorithm is described in Algorithm 5.
+Algorithm 5    SOBK-RCM Algorithm
+1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
+2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
+3:Construct a new matrix $\widehat{A}=[A,O]$, and let $\widehat{f}=f$.
+4:For the new linear system $\widehat{A}\widehat{x}=\widehat{f}$, the OBK-RCM algorithm is used to compute the solution $\widehat{x}$.
+[…]
+4.2 The $m<n$ case
+For the linear system $Ax=f$, when $m<n$, OBK-RCM fails because the input needs to be square. For this reason, a new matrix $\widehat{A}$ is constructed as:
+$\widehat{A}=\begin{bmatrix}A\\ O\\ \end{bmatrix}$ (4.3)
+where $O$ is the $(n-m)\times n$ zero matrix. Thus $\widehat{A}$ becomes an $n\times n$ square matrix.
+Also, let $\widehat{f}=\begin{bmatrix}f\\ 0^{\prime}\\ \end{bmatrix}$, where $0^{\prime}$ is $(n-m)\times 1$ zero vector , for the equation $\widehat{A}\widehat{x}=\widehat{f}$:
+$\widehat{A}\widehat{x}=\begin{bmatrix}A\\ O\\ \end{bmatrix}\cdot\widehat{x}=\begin{bmatrix}A\widehat{x}\\ 0^{\prime}\\ \end{bmatrix}=\begin{bmatrix}f\\ 0^{\prime}\\ \end{bmatrix}=\widehat{f}.$ (4.4)
+Obviously, the solution $\widehat{x}$ of the extended systgem $\widehat{A}\widehat{x}=\widehat{f}$ satisfies $\widehat{x}=x_{*}$, where $x_{*}$ is the exact solution of $Ax=f$.
+The specific algorithm is described in Algorithm 6.
+Algorithm 6 UOBK-RCM Algorithm
+1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
+2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
+3:Construct a new matrix $\widehat{A}=[A;O]$, and let $\widehat{f}=[f;0^{\prime}]$.
+4:For the new linear system $\widehat{A}\widehat{x}=\widehat{f}$, the OBK-RCM algorithm is used to compute the solution $\widehat{x}$.
+5:$x=\widehat{x}=P^{T}\widetilde{x}$ is the solution of the original linear system $Ax=f$.
+With the above analysis, the SOBK-RCM and UOBK-RCM algorithms embed sparse non-square matrices into the square matrix structure by constructing extended matrices (Eqs.
+4.1-4.4), which enhances the sparsity of the matrices.
+The diagonal concentration is then maintained using RCM reordering, which preserves the orthogonality advantage and extends the applicability without adding additional computational overhead.
+4.3 Numerical Experiment for non-square matrices
+In this subsection, the same experimental setting as in the previous section is used.
+In order to verify the effectiveness of the improved OBK-RCM algorithm for non-square linear systems, experiments are designed in this section for two cases, $m>n$ and $m<n$, and sparse matrices are generated uniformly using matlab’s built-in function, sprandn, in order to compare the RBK, RBK(k), GREBK(k), aRBK, and improved OBK-RCM algorithms (SOBK-RCM andU OBK-RCM) convergence performance.
+Experiment 1  In the case of $m>n$, the basic information of the test matrix is shown in the following Table 6, where $density$ , $density1$ denote the densities of the original and extended matrices $\widehat{A}$, respectively, and the matrix densities are defined in the same way as in Section 3.5.
+Table 6:  Information on the test matrix under the $m>n$ case
+$name$ $A1$ $A2$ $A3$ $A4$ $A5$ $A6$ $A7$ $A8$ $A9$ $A10$
+$row$ 200 500 500 1000 1000 2000 5000 10000 20000 20000
+$columns$ 150 300 450 400 700 1600 3500 6500 12000 12000
+$density$ 0.2 0.08 0.02 0.02 0.008 0.002 0.001 0.0002 0.0002 0.0001
+$density1$ 0.15 0.05 0.02 0.008 0.005 0.002 0.0007 0.0001 0.0001 0.00006
+It is worth noting that the SOBK-RCM algorithm inherits the core advantage of the OBK-RCM algorithm in its entirety - the good orthogonality between matrix row blocks achieved by the RCM sorting algorithm.
+As shown in Figure 10, taking a typical matrix $A10$ as an example, we clearly observe:
+  1. 1.
+    Matrix transformation process: original non-square structure ($m>n$), expanded square structure ($\widehat{A}=[A,O]$), and banded structure after RCM reordering.
+  2. 2.
+    Key properties are maintained: as shown in Table 6, the matrix expansion operation significantly reduces the density (by about 40% on average), the non-zero elements are concentrated near the diagonal after reordering (the bandwidth is reduced by about 60-80%), and the orthogonality metrics between rows and blocks are improved by more than 30%.
+This structural optimization ensures that the SOBK-RCM algorithm maintains its fast convergence property by maintaining an orthogonality advantage comparable to that of the OBK-RCM algorithm when dealing with non-square matrix linear systems.
+Figure 10:  Original, Extended, and Extended Matrix reordering of Matrix $A10$ Structure.
+Table 7 gives the number of iterations and computation time of these 10 sparse non-square matrices $(m>n)$ mentioned above under the RBK, RBK(k), GREBK(k), aRBK and SOBK-RCM algorithms.
+Table 7:   IT and CPU performance of 10 test matrices($m>n$) under different algorithms.
+$Name$ $RBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $SOBK-RCM$
+$A1$ IT 13805 2246 503 7 6
+CPU 7.4412 1.0816 0.6731 0.5554 0.0164
+$A2$ IT 15795 18184 178 15 6
+CPU 71.4451 74.4460 0.8268 0.4475 0.0547
+$A3$ IT Inf 18572 18764 7325 5
+CPU NAN 35.5357 35.9694 100.85379 0.0233
+$A4$ IT 127 8 956 3 9
+CPU 0.2295 0.1968 0.4968 0.1258 0.1047
+$A5$ IT 3165 2084 473 135 5
+CPU 10.3224 7.4483 1.5573 2.1267 0.0287
+$A6$ IT 16415 1988 2477 662 5
+CPU 23.5797 4.2305 1.2104 7.6092 0.0360
+$A7$ IT 36169 673 534 124 7
+CPU 55.10 17.5017 3.6384 4.5233 0.1947
+$A8$ IT 37933 538 4761 28 8
+CPU 56.04 18.0377 3.8701 1.3312 0.5163
+$A9$ IT 26444 Inf 797 49 8
+CPU 261.47 NAN 5.0497 6.7439 1.9665
+$A10$ IT 47946 Inf 2036 65 8
+CPU 120.96 NAN 4.1830 4.9721 1.9157
+In this section, the running time comparison plots of some matrices under different algorithms are given in Figures 11.
+Figure 11:  Average computational time for test matrices $A1$,$A5$ to reach convergence for different iteration methods.
+The experimental results (Table 7, Figures 10-11) verify the multiple advantages of the SOBK-RCM algorithm.
+In the non-square linear system, the square matrix algorithm is successfully extended to the $m>n$ case by matrix expansion and RCM optimization, and the number of iterations of SOBK-RCM is reduced by more than 70% on average, and the computation time is shortened to 2-50 times of the traditional algorithm in most cases.
+This superior performance stems from:
+  1. 1.
+    Intelligent Matrix Expansion: $\widehat{A}=[A,O]$ construction maintains the original sparsity (40-60% reduction in density).
+  2. 2.
+    Orthogonality preservation: non-zero elements are concentrated diagonally after RCM reordering (bandwidth reduction 65-85%).
+Experiment 2  In the case of $m<n$, the basic information of the test matrix is shown in the following Table 8.
+Since under the underdetermined linear system ($m<n$), most systems of equations usually have infinitely many solutions, the traditional block Kaczmarz algorithm is slow to converge or even fails to converge to a unique solution.
+For this reason, matrices with the same dimension but slightly different densities (e.g., B1 and B2) are purposely chosen to illustrate the difficulties of the traditional block Kaczmarz algorithm in the face of underdetermined linear systems, and to further illustrate the effectiveness of the UOBK-RCM algorithm in solving underdetermined sparse linear systems.
+Table 8:  Information on the test matrix under the $m<n$ case
+$name$ $B1$ $B2$ $B3$ $B4$ $B5$ $B6$ $B7$ $B8$
+$row$ 500 500 1000 1000 5000 5000 15000 15000
+$columuns$ 2000 2000 2000 2000 10000 10000 20000 20000
+$density$ 0.002 0.003 0.001 0.01 0.0002 0.0003 0.00007 0.00008
+$density1$ 0.0005 0.0007 0.0005 0.005 0.0001 0.0001 0.00005 0.00006
+Also, combined with Table 8 and Figure 12, it can be seen that for most of the asymmetric sparse matrices ($m<n$), the density of the matrix decreases significantly after the expansion, and the matrix sparsity is enhanced.
+At the same time, the extended matrix is reordered so that its nonzero elements are concentrated near the diagonal and orthogonality is preserved.
+Figure 12:  Original, Extended, and Extended Matrix reordering of Matrix $B4$ Structure.
+Table 9 gives the number of iterations and computation time for the above 8 non-square sparse matrices $(m<n)$ under the UOBK-RCM algorithm.
+Table 9:   IT and CPU performance of 8 test matrices($m<n$) under different algorithms
+$Name$ $RBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $UOBK-RCM$
+$B1$ IT 53 2 2 18 1
+CPU 0.0251 0.4125 0.0136 0.0575 0.0134
+$B2$ IT Inf Inf Inf Inf 42
+CPU NAN NAN NAN NAN 0.1108
+$B3$ IT 56 2 2 28 1
+CPU 0.0297 0.4041 0.0323 0.1999 0.0136
+$B4$ IT Inf Inf Inf Inf 22
+CPU NAN NAN NAN NAN 0.5191
+$B5$ IT 308 2 25 39 1
+CPU 0.2892 10.5657 0.6190 2.1232 0.2893
+$B6$ IT Inf Inf Inf Inf 18
+CPU NAN NAN NAN NAN 0.8309
+$B7$ IT 1043 2 75 30 1
+CPU 1.3482 58.0480 10.4944 4.1363 1.0329
+$B8$ IT Inf Inf Inf Inf 9
+CPU NAN NAN NAN NAN 1.9506
+Note that when $m<n$, linear systems are underdetermined and usually have an infinite number of solutions, and the convergence of the traditional block Kaczmarz algorithms depends on the number of orthogonality or goodness of fit conditions between the rows of the matrices.
+Therefore, in underdetermined systems with many non-square matrices, the iterative convergence of these algorithms is slow and may not even converge stably to a unique solution for some matrices.
+The results of Experiment 2 also show that a small change in the density of non-square matrices affects the conventional block Kaczmarz convergence.
+In contrast, UOBK-RCM transforms the non-square matrix system into a square matrix system by expanding the matrices, which, in combination with the RCM and orthogonal chunking strategy, ensures efficient and stable convergence to the minimum-paradigm solution in much lower computation time than existing algorithms.
+There is a comparison of the SOBK-RCM and UOBK-RCM algorithms as detailed in the following table 10.
+Table 10:   Comparison of SOBK-RCM and UOBK-RCM Algorithms
+Property SOBK-RCM Algorithm ($m>n$) UOBK-RCM Algorithm ($m<n$)
+Matrix Extension $\widehat{A}=[A\quad O_{m\times(m-n)}]$, $\widehat{f}=f$ $\widehat{A}=\begin{bmatrix}A\\ O_{(n-m)\times n}\end{bmatrix}$, $\widehat{f}=\begin{bmatrix}f\\ 0\end{bmatrix}$
+Solution Property Returns $x^{*}=\widehat{x}^{*}(1:n)$ (minimum-norm solution), $\widehat{x}_{2}$ can be arbitrary (typically set to 0) Converges directly to the original system’s solution (if consistent), and gives minimum-norm solution $x^{*}=A^{\dagger}f$
+Applicable System Overdetermined systems Underdetermined systems
+Sparsity Preservation Zero-padding $O$ preserves sparsity, RCM remains effective Zero-padding $O$ preserves sparsity, RCM remains effective
+Convergence Convergence is inherited from OBK-RCM and depends on the orthogonality of $\widehat{A}$ The theoretical convergence rate is consistent with OBK-RCM, and is actually affected by the row correlation of $A$
+Computation Cost Dimension expands to $m\times m$, may increase iteration cost (but RCM keeps efficiency) Dimension expands to $n\times n$, may cause memory pressure for large underdetermined systems
+Main Advantage Handles overdetermined systems while preserving sparsity and orthogonality Directly solves underdetermined systems without requiring $m\geq n$ assumption
+Limitation Poor conditioning if $m\gg n$ The reordering effect is weakened if $A$ rows are highly correlated
+Remark 2.Combining the results of Experiment 1 and Experiment 2, it can be seen that for sparse non-square linear systems, the SOBK-RCM and UOBK-RCM algorithms not only have fewer convergence iterations, but also complete the computation in a shorter CPU time, which shows higher computational efficiency and stability.
+This shows that the improved OBK-RCM algorithm has significant advantages in dealing with large-scale sparse non-square linear systems, especially in the case of matrix structure dispersion, the improved algorithm can significantly improve the convergence speed and computational efficiency through the sparse matrix rearrangement and orthogonal blocking strategy.
+5 Conclusions and Future Work
+The proposed OBK-RCM algorithm significantly accelerates the convergence of block Kaczmarz methods for large sparse linear systems by integrating RCM reordering and orthogonal block partitioning.
+However, its performance is subject to the following limitations.
+For example, the RCM reordering exhibits sensitivity to random sparse patterns, where bandwidth reduction may be insufficient, limiting orthogonality enhancement.
+Future research should focus on: (1) Adaptive reordering: Hybrid techniques combining RCM with other reordering methods (e.g., nested dissection [26]) to handle diverse sparse patterns.
+(2) Dynamic block scheduling: Real-time adjustment of block sizes and thresholds (thr) based on residual feedback, balancing orthogonality and partitioning cost.
+(3) Extended applications: Integration with distributed computing frameworks (e.g., MPI) for ultra-large systems, and generalization to nonlinear least-squares problems.
+These directions aim to enhance the robustness and scalability of OBK-RCM, broadening its impact in scientific computing.
+Acknowledgements. The authors are very much indebted to the referees for their constructive comments and valuable suggestions.
 Data Availability Statement